### v1

# On the existence of forward self-similar solutions to the two-dimensional Navier-Stokes equations

Abstract.
We establish the global existence of forward self-similar solutions to the two-dimensional incompressible Navier-Stokes equations for any divergence-free initial velocity $u_{0}$ that is homogeneous of degree 1-1 and locally Hölder continuous. This result requires no smallness assumption on the initial data. In sharp contrast to the three-dimensional case, where (1)(-1)-homogeneous vector fields are locally square-integrable, the 2D problem is critical in the sense that the initial kinetic energy is locally infinite at the origin, and the initial vorticity fails to be locally integrable. Consequently, the classical local energy estimates are not available. To overcome this, we decompose the solution into a linear part solving the heat equation and a finite-energy perturbation part. By exploiting a kind of inherent cancellation relation between the linear part and the perturbation part, we can control interaction terms and establish the $H^{1}$-estimates for the perturbation part. Further investigating the corresponding Leray system in weighted Sobolev space, we can derive an optimal pointwise estimate. This gives the faster decay of the perturbation part at infinity and enables us to construct global-in-time self-similar solutions.
Key words and phrases:
Navier-Stokes Equation, forward Self-Similar Solution, Large initial data, two-dimensional
2010 Mathematics Subject Classification:
35B40, 35C06, 35Q30, 76D05
1. Introduction and Main Results
We consider the two-dimensional Navier-Stokes equations
[…]
with initial condition
(1.2) $u(x,0)=u_{0}(x).$
Here $u(x,t):\mathbb{R}^{2}\times[0,+\infty)\to\mathbb{R}^{2}$ is the velocity vector field and the scalar $\pi(x,t):\mathbb{R}^{2}\times[0,+\infty)\to\mathbb{R}$ is the pressure. The problem (1.1) is invariant under the scaling
$u(x,t)\to u_{\lambda}(x,t)=\lambda u(\lambda x,\lambda^{2}t),$
$\pi(x,t)\to\pi_{\lambda}(x,t)=\lambda^{2}\pi(\lambda x,\lambda^{2}t).$
$u_{0}(x)\to u_{0,\lambda}(x)=\lambda u_{0}(\lambda x).$
In particular, the solution $u$ is called self-similar or scale-invariant if $u_{\lambda}=u$ and $\pi_{\lambda}=\pi$ for any $\lambda>0$. Similarly, we say that the initial condition $u_{0}$ is self-similar or scale-invariant, if $u_{0}$ satisfies $\lambda u_{0}(\lambda x)=u_{0}(x)$ any for $\lambda>0$.
The study of self-similar solutions is motivated by their fundamental roles in understanding the structure of singularities and the asymptotic behavior of the Navier-Stokes equations. In the context of the regularity problem, (backward) self-similar solutions often serve as natural candidates for blow-up profiles. Although the 2D Navier-Stokes equations are well-known to be globally well-posed for smooth data with finite energy, the forward self-similar solutions with singular initial data provide a critical setting to test the robustness of the theory.
Furthermore, self-similar solutions are quite useful for describing the long-time asymptotics of general solutions, which themselves usually evolve from the singular initial data. For the the Navier-Stokes equations in 2D, it is known that a special class of self-similar solutions called Oseen vortices dominate the long-time behaviour of general solutions to (1.1) with integrable initial vorticity [14, 15, 11]; see also the discussion below. We mention that for the 2D Navier-Stokes equations with integrable vorticity, the solution can have infinite energy in general.
Our main goal is to construct a global self-similar solution for any self-similar initial velocity field, generally with non-integrable vorticity. Before stating our result, we give the definition of the solution we will use; this definition is essentially the same as [26, Definition 8.18].
Definition 1.1.
A vector field $u$ defined on $\mathbb{R}^{2}\times[0,\infty)$ is called an energy perturbed solution with divergence-free initial data $u_{0}\in L^{2,\infty}(\mathbb{R}^{2})$ if it satisfies (1.1) in the sense of distributions, and
  • (i)
    $u-e^{t\Delta}u_{0}\in L^{\infty}(0,T;L^{2}(\mathbb{R}^{2}))\cap L^{2,\infty}(0,T;H^{1}(\mathbb{R}^{2}))$ for any $T>0$,
  • (ii)
    $\lim_{t\to 0}\|u(t)-e^{t\Delta}u_{0}(t)\|_{L^{2}(\mathbb{R}^{2})}=0$,
where
$(e^{t\Delta}u_{0})(x,t)=\int_{\mathbb{R}^{2}}\frac{1}{4\pi t}e^{-\frac{|x-z|^{2}}{4t}}u_{0}(z)dz.$
is the solution to the heat equation with the same initial data $u_{0}$.
The main result in this paper is the following theorem.
Theorem 1.1.
Assume $u_{0}(x)\in C^{0,\beta}_{loc}$ for some $0<\beta\leq 1$ is a divergence-free, self-similar initial data in $\mathbb{R}^{2}\setminus\{0\}$, and that $\int_{\partial S^{1}}u_{0}\cdot nd\sigma=0$. Suppose that
$A=\|u_{0}\|_{C^{0,\beta}(S^{1})}<+\infty,$
here $S^{1}$ is the unit circle on $\mathbb{R}^{2}$. Then there is a global smooth self-similar solution $u(x,t)\in C^{\infty}(\mathbb{R}^{2}\times(0,\infty))$ to the Cauchy problem (1.1)-(1.2). Moreover, the solution $u$ is an energy perturbed solution in the sense of Definition 1.1. Finally, we have the following pointwise estimates
$|u(x,t)|\leq\frac{C(A,\beta)}{|x|+\sqrt{t}},\quad~|u(x,t)-e^{t\Delta}u_{0}|\leq C(A,\beta)\frac{\sqrt{t}}{|x|^{2}+t},$
and
$|\nabla u(x,t)-\nabla e^{t\Delta}u_{0}|\leq C(A,\beta)\frac{\sqrt{t}^{\beta}}{(|x|+\sqrt{t})^{2+\beta}}.$
If, in addition, the initial data $u_{0}$ belongs to $C^{2}(S^{1})$, then we have the following optimal decay rate estimates
(1.3) $|u(x,t)-e^{t\Delta}u_{0}|\leq C(\|u_{0}\|_{C^{2}(S^{1})})\frac{t}{(|x|+\sqrt{t})^{3}}\ln\left(1+\frac{|x|}{\sqrt{t}}\right).$
Here, $C^{0,\beta}_{loc}$ is the standard local Hölder space and when $\beta=1$, $C^{0,1}_{loc}$ is the local Lipshitz space.
Remark 1.1.
The decay rate (1.3) is optimal and cannot be improved in general, even if $u_{0}$ has higher regularity. One can see the precise asymptotic expansions in [8] when the initial data is small. The remainder term $\frac{t}{(|x|+\sqrt{t})^{3}}\ln\left(1+\frac{|x|}{\sqrt{t}}\right)$ can not vanish unless the orthogonality relations
$\int_{S^{1}}u_{0,1}^{2}=\int_{S^{1}}u_{0,2}^{2}\textrm{ and }\int_{S^{1}}u_{0,1}u_{0,2}=0$
are satisfied.
Remark 1.2.
It seems that the condition $\int_{\partial B_{1}}u_{0}\cdot nd\sigma=0$ is necessary for Theorem 1.1 in view of the Jeffery–Hamel solutions [17], which are steady self-similar solutions with generally non-zero flux, and thus are always singular at the origin. We mention that this is different with the 3D case, where any divergence-free ($L^{\infty}_{loc}$) similar vector field $u_{0}$ on $\mathbb{R}^{3}\setminus\{0\}$ automatically satisfies ${\rm div\thinspace}u_{0}=0$ across the origin.
Remark 1.3.
It is possible to reduce the regularity assumptions in Theorem 1.1 to construct large self-similar solutions. See [12, 7, 6, 1] for the results of the 3D case with rough initial data.
Remark 1.4.
We recently got to know that Dallas Albritton, Julien Guillod, Mikhail Korobkov, and Xiao Ren ([24]) obtained a result similar to Theorem 1.1, where the solution at $t=1$ decays as $|x|^{-1-\beta}$ if the initial data $u_{0}$ belongs to $C^{0,\beta}(S^{1})$.
The existence of solutions in a scale-invariant space has been studied extensively, including $L^{n}$, $L^{n,\infty}$, Besov spaces and $BMO^{-1}$ in $n$ dimensions (usually $n=2,3$), typically under smallness assumptions, see [3, 10, 16, 19]. Especially, the small data existence of forward self-similar solutions follows from the uniqueness results in these studies. One may also refer to [26, Chapter 8] and the reference therein for more references.
For $\mathbb{R}^{3}$, recently, the seminal work of Jia and Šverák [18] established the global existence of smooth self-similar solutions for the 3D Navier-Stokes equations with any Hölder continuous self-similar initial data on $\mathbb{R}^{3}\setminus\{0\}$. The main techniques of their approach are to derive the local-in-space Hölder estimates based on the $L_{loc}^{2}$-energy estimates established by Lemarié-Rieusset for the local Leray solutions in [23]. Later in [20], the self-similar solutions were constructed based on the a priori $H^{1}$-estimate obtained for the equivalent Leray equations by Leray’s method of contradiction, which is a kind of blow-up argument. In [4, 5], Bradshaw and Tsai gave a rather direct method to establish the $H^{1}$-estimates for both discretely self-similar solutions and rotated discretely self-similar solutions, and then constructed solutions via the Galerkin approximation. For more related results and constructions of self-similar solutions for the Navier-Stokes equations with more general self-similar initial data, for example, for initial data that lies in $L^{2}_{loc}(\mathbb{R}^{3})$ and Besov space, one can refer to [25, 27, 12, 7, 6, 1]. For the constructions of self-similar solutions for the fractional Navier-Stokes equations and other related models, such as MHD and SQG equation, one can refer to [22, 21, 29, 28, 2].
For $\mathbb{R}^{2}$, note that the (-1)-homogeneous divergence-free velocity field $u_{0}(x)$ has constant circulation on any circle $\partial B_{R}$, hence can also be decomposed as
$u_{0}(x)=\frac{\alpha}{2\pi}\frac{x^{\perp}}{|x|^{2}}+\tilde{u}_{0},$
with $\tilde{u}_{0}$ has zero circulation. Taking curl, we have the corresponding vorticity
$\omega_{0}(x)=\alpha\delta+\textrm{curl }\tilde{u}_{0}.$
When $\tilde{u}_{0}=0$, an explicit class of self-similar solutions taking $\alpha\delta$ as initial data (in vorticity formulation of the Navier-Stokes equations) exists for any $\alpha\in\mathbb{R}$ without smallness restrictions. They are known as Oseen vortices, and the velocity and vorticity are given by:
$u(x,t)=\frac{\alpha}{2\pi}\frac{x^{\perp}}{|x|^{2}}(1-e^{-\frac{|x|^{2}}{4t}}),~\omega(x,t)=\frac{\alpha}{4\pi t}e^{-\frac{|x|^{2}}{4t}}.$
Indeed the vorticity $\omega(x,t)=\frac{\alpha}{4\pi t}e^{-\frac{|x|^{2}}{4t}}$ just solves the heat equation in this case. Oseen vortices characterize the long-time asymptotic behavior of general solutions to (1.1) with integrable initial vorticity [14, 15, 11]. On the other hand, there is no previous known existence result of self-similar solutions for large $\tilde{u}_{0}$, even for the special case $\alpha=0$.
The main difficulty in proving Theorem 1.1, compared to the 3D case, lies in the non-integrability of the initial data, even locally, at the level of both velocity and vorticity on $\mathbb{R}^{2}$. In 2D, a self-similar initial velocity $|u_{0}(x)|\sim|x|^{-1}$ does not belong to $L_{loc}^{2}(\mathbb{R}^{2})$, hence, the local Leray solutions developed in Lemarié-Rieusset in [23] is not applicable in 2D. On the other hand, the initial vorticity does not belong to $L_{loc}^{1}(\mathbb{R}^{2})$ and even can not be viewed as a finite measure as $|\textrm{curl}~u_{0}(x)|\sim|x|^{-2}$. Consequently, we lack a foundational quantity to initiate the estimates at both the velocity and the vorticity level.
To overcome this lack of initial integrability, we adopt a decomposition strategy. We postulate that the singular behavior of the solution is dominated by the linear evolution. Let $e^{t\Delta}u_{0}$ be the caloric lift of the initial data, i.e., the solution to the heat equation starting from the same initial data $u_{0}$. While $e^{t\Delta}u_{0}$ retains the non-$L^{2}$ character of the initial data, it is explicitly computable and smooth for $t>0$. The unknown remainder term $u_{\text{re}}=u-e^{t\Delta}u_{0}$ is then expected to possess better decay properties. By substituting this decomposition into the Navier-Stokes equations and using the self-similarity variables, we derive a perturbed Leray system. The core of our proof lies in establishing global-in-time energy estimates for $u_{\text{re}}$, depending only on initial data. Due to the self-similarity, we can focus on the energy estimates at $t=1$ only.
Our key observation is that there is some inherent cancellation relation between $e^{\Delta}u_{0}$ and $u_{\text{re}}(\cdot,1)$. Using this relation, we derive a crucial identity, see (2.10) below. This identity implies the Dirichlet energy of $u_{\text{re}}(\cdot,1)$ can be controlled by that of the linear part $e^{\Delta}u_{0}$, which is finite by using explicit estimates for $e^{\Delta}u_{0}$. Starting from this Dirichlet energy estimate, we can obtain the $L^{2}$ control for the $u_{\text{re}}(\cdot,1)$ itself via a suitable choice of the multiplier and the interpolation inequality. One can refer to Section 2.2 for a more detailed outline of the proof.
Once we get initial energy estimates for $u_{\text{re}}(\cdot,1)$, we use weighted estimates inspired by [22] to derive pointwise estimate for $u_{\text{re}}(\cdot,1)$. Specifically, we choose appropriate test functions to derive the following key estimate
$\sqrt{1+|x|^{2}}u_{\text{re}}(x,1)\in H^{2}(\mathbb{R}^{2}).$
Based on this regularity and Sobolev embedding, we have the following pointwise estimate of $u_{\text{re}}$ for $x$,
$|u_{\text{re}}(x,1)|\leq\frac{C}{1+|x|}\textrm{ for all }x\in\mathbb{R}^{2}.$
With this decay estimate in hand, we can eventually improve the decay estimates by using the linear theory of Stokes and the heat equation to have that
$|u_{\text{re}}(x,1)|\leq\frac{C}{1+|x|^{2}}\textrm{ for all }x\in\mathbb{R}^{2}.$
This faster decay compared with the linear part $e^{\Delta}u_{0}$, together with standard interior regularity, gives compactness, so that we can employ the Leray-Schauder degree theory to find a self-similar solution to the Navier-Stokes equation (1.1). On the other hand, if $u_{0}\in C^{2}_{loc}(\mathbb{R}^{2})$, we can utilize the computations in [8] to show the optimal decay rate (1.3).
The rest of the paper is organized as follows. We give the proof of the energy estimates for the remainder term $u_{\text{re}}(\cdot,1)=u(\cdot,1)-e^{\Delta}u_{0}$ in Section 2. Next, we derive pointwise decay estimates for $u_{\text{re}}$ by establishing the weighted energy estimates in Section 3. Finally, we prove the main theorem in Section 4 by using the Leray-Schauder degree theory.
2. The Leray equations and the Energy estimates
The primary objective of this section is to establish the $H^{1}(\mathbb{R}^{2})$- energy estimates for the remainder term defined by
$u_{\text{re}}(x,1)=\left.u-e^{t\Delta}u_{0}\right.|_{t=1}=u(x,1)-e^{\Delta}u_{0}(x).$
Furthermore, we demonstrate that $u_{\text{re}}(x,1)$ is $L^{p}$-integrable for any $p\in(1,+\infty)$. This implies that $u_{\textrm{re}}$ decays like $\frac{1}{1+|x|^{2}}$ in an average sense. The local regularity of $u_{\text{re}}(x,1)$ and $e^{\Delta}u_{0}$ is not a big issue, indeed they are smooth due to the parabolic smoothing effect, once we have some initial local integrability. However, they decay slowly at spatial infinity, and the key point is to get the global integrability to initiate the subsequent analysis. Before stating the main result in this section, we first introduce self-similar variables and the Leray equations.
2.1. Self-similar variables and the Leray equations
Our analysis will be conducted in self-similar variables:
(2.1) $y=\frac{x}{\sqrt{t}},~s=\log(t).$
We define $v(y,s)$ as
(2.2) $u(x,t)=\frac{1}{\sqrt{t}}v(y,s)=\frac{1}{\sqrt{t}}v\left(\frac{x}{\sqrt{t}},\log t\right).$
It follows that if $u$ solves (1.1), then $v$ satisfies the time-dependent Leray equations:
(2.3) $\left\{\begin{aligned} &\frac{\partial v}{\partial s}-\Delta v-\frac{1}{2}v-\frac{1}{2}y\cdot\nabla v+v\cdot\nabla v+\nabla q=0,\\ &\textrm{ div }v=0.\end{aligned}\right$.
The self-similarity of $u$ implies that $v$ is independent of $s$, and in this case $v(y)=u(y,1)$. We have the following stationary Leray equations for $v$
(2.4) $\left\{\begin{aligned} &-\Delta v-\frac{1}{2}v-\frac{1}{2}y\cdot\nabla v+v\cdot\nabla v+\nabla q=0,\\ &\textrm{ div }v=0.\end{aligned}\right$.
As is expected $v\sim\frac{C}{1+|y|}$, hence, $v\notin L^{2}(\mathbb{R}^{2})$. To circumvent this, we subtract the linear part from $v$ using the following decomposition. Let $e^{t\Delta}u_{0}$ be the solution to the heat equation with initial data $u_{0}$, that is
$e^{t\Delta}u_{0}(x,t)=\int_{\mathbb{R}^{2}}\frac{1}{4\pi t}e^{-\frac{|x-z|^{2}}{4t}}u_{0}(z)dz.$
Direct calculation shows that $e^{t\Delta}u_{0}$ inherits the self-similarity of $u_{0}$:
$\lambda(e^{t\Delta}u_{0})(\lambda x,\lambda^{2}t)$ $=\lambda\int_{\mathbb{R}^{2}}\frac{1}{4\pi\lambda^{2}t}e^{-\frac{|\lambda x-z|^{2}}{4\lambda^{2}t}}u_{0}(z)dz$
$=\lambda\int_{\mathbb{R}^{2}}\frac{1}{4\pi\lambda^{2}t}e^{-\frac{|\lambda x-\lambda z|^{2}}{4\lambda^{2}t}}u_{0}(\lambda z)\lambda^{2}dz$
$=\int_{\mathbb{R}^{2}}\frac{1}{4\pi t}e^{-\frac{|x-z|^{2}}{4t}}u_{0}(z)dz=e^{t\Delta}u_{0}(x,t),$
where in the last line we have used that $u_{0}$ is self-similar. We define $v_{0}(y)$ by
$v_{0}(y)=\sqrt{t}(e^{t\Delta}u_{0})(x,t),$
then obviously we have
(2.5) $v_{0}(y)=(e^{t\Delta}u_{0})\left(\frac{x}{\sqrt{t}},1\right)=(e^{\Delta}u_{0})(y).$
Let $v_{\text{re}}$ be the remainder part of $v$ when subtracting the linear part $v_{0}$, that is
(2.6) $v_{\text{re}}(y)=v(y)-v_{0}(y)=u_{\text{re}}(y,1).$
Our primary goal will then be to get a priori estimates for $v_{\text{re}}(y)$. It is easy to see that $v_{0}$ satisfies
(2.7) $\Delta v_{0}-\frac{1}{2}v_{0}-\frac{1}{2}y\cdot\nabla v_{0}=0,~{\rm div\thinspace}v_{0}=0.$
Hence, $v_{\text{re}}$ satisfies
(2.8) $\left\{\begin{aligned} &-\Delta v_{\text{re}}-\frac{1}{2}v_{\text{re}}-\frac{1}{2}y\cdot\nabla v_{\text{re}}+v_{0}\cdot\nabla v_{\text{re}}+v_{\text{re}}\cdot\nabla v_{0}+v_{\text{re}}\cdot\nabla v_{\text{re}}+\nabla q=-v_{0}\cdot\nabla v_{0},\\ &{\rm div\thinspace}v_{\text{re}}=0.\end{aligned}\right$.
The main objective of this Section is to prove the following.
Theorem 2.1.
Let the divergence-free, self-similar initial data $u_{0}(x)\in C^{0,\beta}_{loc}$ be given, $0<\beta\leq 1$. We denote
$A=\|u_{0}\|_{C^{0,\beta}(S^{1})}<+\infty.$
Assume $u$ is a self-similar solution to (1.1) in the sense of Definition 1.1. Consider the corresponding solution to (2.8) in which $v_{0}=e^{\Delta}u_{0}$, then there exists a constant $C=C(A,\beta)$ depending only on $A$ and $\beta$, such that
$\int_{\mathbb{R}^{2}}|v_{\text{re}}|^{2}+|\nabla v_{\text{re}}|^{2}dx\leq C(A,\beta).$
Moreover, we have that for any positive constant $1<p\leq\infty$, there is a constant $C=C(A,\beta,p)$ depending on $A$, $p$, and $\beta$ that
$\|v_{\text{re}}\|_{L^{p}(\mathbb{R}^{2})}\leq C(A,\beta,p).$
Furthermore, we have the following estimates for the pressure $q$ in (2.8). For any positive constant $1<p<\infty$, we have
$\|q\|_{L^{p}(\mathbb{R}^{2})}\leq C(A,\beta,p).$
and also
$\|\nabla q\|_{L^{2}(\mathbb{R}^{2})}\leq C(A,\beta).$
2.2. Main difficulties, key ideas and the outline for the energy estimates
In this section, we give the main idea for the a priori estimates of solutions in Theorem 2.1.
Formally multiplying (2.8) by $v_{\text{re}}$ and integrating by parts, we arrive at the energy identity
(2.9) $\int_{\mathbb{R}^{2}}|\nabla v_{\text{re}}|^{2}+(v_{\text{re}}\cdot\nabla v_{0})\cdot v_{\text{re}}=-\int_{\mathbb{R}^{2}}(v_{0}\cdot\nabla v_{0})\cdot v_{\text{re}}.$
The main difficult lies in controlling the term $\int_{\mathbb{R}^{2}}(v_{\text{re}}\cdot\nabla v_{0})\cdot v_{\text{re}}$. If one uses integration by parts, we have the following estimates
$\int_{\mathbb{R}^{2}}(v_{\text{re}}\cdot\nabla v_{0})\cdot v_{\text{re}}=-\int_{\mathbb{R}^{2}}(v_{\text{re}}\cdot\nabla v_{\text{re}})\cdot v_{0}\leq\|v_{0}\|_{L^{\infty}}\int_{\mathbb{R}^{2}}(|v_{\text{re}}|^{2}+|\nabla v_{\text{re}}|^{2}),$
yet $\|v_{\text{re}}\|_{L^{2}(\mathbb{R}^{2})}$ cannot be controlled at this stage. This situation differs from the three-dimensional case, where there is an additional term $\frac{1}{4}\int_{\mathbb{R}^{2}}|v_{\text{re}}|^{2}$ that appears in the left-hand side of (2.9) and help to close the energy estimates, provided $\|v_{0}\|_{L^{\infty}}$ is small via a suitable cut-off; see [4, 5].
Our key insight is that an inherent structural relationship exists between $v_{0}$ and $v_{\text{re}}$, which can be more effectively utilized than a simple size estimate of $\int(v_{\text{re}}\cdot\nabla v_{0})\cdot v_{\text{re}}$. Our way to exploit this inherent relation is as follows. We multiply (2.8) by $(v_{0}+v_{\text{re}})$ and multiply (2.7) by $v_{\text{re}}$. Summing the resulting equations and integrating over $\mathbb{R}^{2}$, we obtain, after integration by parts, a very crucial identity
(2.10) $\int_{\mathbb{R}^{2}}|\nabla v_{\text{re}}|^{2}dy+2\int_{\mathbb{R}^{2}}\nabla v_{0}\cdot\nabla v_{\text{re}}dy=0.$
This, together with the Hölder inequality, leads to
(2.11) $\int_{\mathbb{R}^{2}}|\nabla v_{\text{re}}|^{2}dy\leq 4\int_{\mathbb{R}^{2}}|\nabla v_{0}|^{2}dy\leq C.$
as $\nabla v_{0}\sim\frac{C}{1+|y|^{1+\beta}}\in L^{2}(\mathbb{R}^{2})$. This provides the initial Dirichlet estimate for $v_{\text{re}}$.
To obtain the $L^{2}$ bound for $v_{\text{re}}$, we return to (2.8) and try to utilize the good term $\frac{1}{2}y\cdot\nabla v_{\text{re}}$ as much as possible. The main idea is to use the multiplier $|v_{\text{re}}|^{p-2}v_{\text{re}}$ for some $1<p<2$. If we formally test (2.8) with $|v_{\text{re}}|^{p-2}v_{\text{re}}$, we have
$\int_{\mathbb{R}^{2}}(p-1)|\nabla v_{\text{re}}|^{2}|v_{\text{re}}|^{p-2}+\left(\frac{1}{p}-\frac{1}{2}\right)|v_{\text{re}}|^{p}$
$\leq$ $-\int_{\mathbb{R}^{2}}(v_{\text{re}}\cdot\nabla v_{0})\cdot|v_{\text{re}}|^{p-2}v_{\text{re}}+\nabla q\cdot|v_{\text{re}}|^{p-2}v_{\text{re}}+(v_{0}\cdot\nabla v_{0})\cdot|v_{\text{re}}|^{p-2}v_{\text{re}}.$
The left hand side contains an additional term $\int_{\mathbb{R}^{2}}\left(\frac{1}{p}-\frac{1}{2}\right)|v_{\text{re}}|^{p}$, which enables us to close the estimates for $\|v_{\text{re}}\|_{L^{p}}$ for any $1<p<2$ eventually. The key issue is to control $\int_{\mathbb{R}^{2}}(v_{\text{re}}\cdot\nabla v_{0})\cdot|v_{\text{re}}|^{p-2}v_{\text{re}}$ and $\int_{\mathbb{R}^{2}}\nabla q\cdot|v_{\text{re}}|^{p-2}v_{\text{re}}$ as the last term on the right hand side is more manageable. To achieve this, we introduce a modified profile of $v_{0}$ to ensure it is small near the origin. Finally, once the $\|v_{\text{re}}\|_{L^{p}}$ estimate is established, the Gagliardo–Nirenberg interpolation inequality combined with (2.11) yields the desired $L^{2}$-boundedness of $v_{\text{re}}$. The higher integrability estimate of $\|v_{\text{re}}\|_{L^{p}}$ when $p>2$ will follow from regularity estimates for the Stokes equation and the bootstrapping argument.
2.3. Proof of the energy estimates
We need to do some preliminary work. We begin with the following decay estimates for the solution to the heat equation. It is more or less standard, and we leave the proof to Appendix A.
Lemma 2.1.
Let $u_{0}$ be divergence-free on $\mathbb{R}^{2}$ and self-similar, and let $0<\beta\leq 1$, denote
$A=\|u_{0}\|_{C^{0,\beta}(S^{1})}<+\infty.$
Then $v_{0}=e^{\Delta}u_{0}(y)$ defined in (2.5) is smooth and divergence-free.
(2.12) $|v_{0}(y)|\leq\frac{C(A)}{1+|y|},~|\nabla^{k}v_{0}(y)|\leq\frac{C(A,\beta,k)}{1+|y|^{1+\beta}},~k\in\mathbb{N},k\geq 1.$
We now proceed to establish energy estimates for $v_{\text{re}}=v-v_{0}$. As discussed in Section 2.2, we need to construct a modified profile of $v_{0}$, which we denote by $v_{1}$, that is globally small via a suitable cut-off procedure. To do so, let $\eta(x)=\eta(|x|)$ be a smooth non-negative cut-off function satisfying
$0\leq\eta(|x|)\leq 1,~\eta(|x|)=1\textrm{ for }|x|\geq 2\textrm{ and }\eta(|x|)=0\textrm{ for }|x|\leq 1.$
We also denote for $R_{0}>1$ that
(2.13) $\eta_{R_{0}}(x)=\eta\left(\frac{x}{{R_{0}}}\right),$
so that there exists some universal constant $C$ that
$|\nabla^{k}\eta_{R_{0}}|\leq\frac{C}{R_{0}^{k}},~\forall k\in\mathbb{N}.$
We define a modified version of $v_{0}$ by
(2.14) $v_{1}=\eta_{R_{0}}v_{0}+w,$
where $w$ is used to make $v_{1}$ is divergence-free, so $w$ solves
${\rm div\thinspace}w=-v_{0}\cdot\nabla\eta_{R_{0}}.$
Explicitly, we can define
(2.15) $w(y)$ $=-\frac{1}{2\pi}\nabla_{y}\int_{\mathbb{R}^{2}}(\ln|y-z|)v_{0}\cdot\nabla\eta_{R_{0}}(z)dz$
$=-\frac{1}{2\pi}\int_{\mathbb{R}^{2}}\frac{y-z}{|y-z|^{2}}v_{0}\cdot\nabla\eta_{R_{0}}(z)dz.$
Remark 2.1.
Such a correction term $w$ is obviously not unique. Typically, it can be constructed with compact support in the annulus $B_{2R_{0}}\setminus B_{R_{0}}$ using the Bogovskii formula, given that $v_{0}\cdot\nabla\eta_{R_{0}}$ is smooth and supported in that region (cf. [13, Theorem III.3.3]). The representation in (2.15), however, may not have compact support.
We now give the estimates for the modified profile $v_{1}=\eta_{R_{0}}v_{0}+w$.
Lemma 2.2 (Estimates for $v_{1}$).
Let $v_{0}$ satisfy the estimates in Lemma 2.1. The vector field $v_{1}$ defined in (2.14) is smooth, divergence-free, and we have the following decay estimates for $v_{1}$ and $w$:
(2.16) $|v_{1}(y)|\leq\frac{C(A,R_{0})}{1+|y|}\textrm{ and }|\nabla^{k}v_{1}(y)|\leq\frac{C(A,\beta,k,R_{0})}{1+|y|^{1+\beta}},~k\in\mathbb{N}^{+},$
and
(2.17) $|\nabla^{k}w(y)|\leq\frac{C(A,\beta,k,R_{0})}{1+|y|^{k+2}},~k=0,1,2\cdots,$
where the constant $C(A,\beta,R_{0})$ depends only on $A$, $\beta$, $k$ and $R_{0}$. Moreover, for any $\epsilon>0$, we can choose $R_{0}=R_{0}(A,\beta,\epsilon)$ in (2.13) large enough so that
(2.18) $\|v_{1}\|_{L^{\infty}(\mathbb{R}^{2})}\leq\epsilon,~\|\nabla v_{1}\|_{L^{\infty}(\mathbb{R}^{2})}\leq\epsilon.$
Proof.
First of all, we have
$|\eta_{R_{0}}v_{0}|\leq\frac{C(A)}{1+|y|},~|\nabla(\eta_{R_{0}}v_{0})|\leq\frac{C(A,\beta)}{1+|y|^{1+\beta}}$
due to Lemma 2.1. Hence, to prove (2.16) and (2.17), it suffices to prove (2.17). We only consider large $y$, as the smoothness of $w$ is easy to see using the smoothness of $v_{0}$. We can assume that $|y|\geq 3R_{0}$. Since $v_{0}\cdot\nabla\eta_{R_{0}}$ is smooth and has compact support, we have
$\int_{\mathbb{R}^{2}}v_{0}\cdot\nabla\eta_{R_{0}}=\int_{B_{2R_{0}}\setminus B_{R_{0}}}\textrm{ div }(\eta_{R_{0}}v_{0})=\int_{\partial B_{2R_{0}}}v_{0}\cdot nd\sigma=0.$
It follows that
$w(y)$ $=-\frac{1}{2\pi}\int_{\mathbb{R}^{2}}\left(\frac{y-z}{|y-z|^{2}}-\frac{y}{|y|^{2}}\right)v_{0}\cdot\nabla\eta_{R_{0}}(z)dz$
$=-\frac{1}{2\pi}\int_{B_{2R_{0}}}\left(\frac{y-z}{|y-z|^{2}}-\frac{y}{|y|^{2}}\right)v_{0}\cdot\nabla\eta_{R_{0}}(z)dz.$
We can rewrite the above as
$w(y)$ $=-\frac{1}{2\pi}\int_{B_{2R_{0}}}\left(\frac{y-z}{|y-z|^{2}}-\frac{y}{|y|^{2}}\right)v_{0}\cdot\nabla\eta_{R_{0}}(z)dz$
$=-\frac{1}{2\pi}\int_{B_{2R_{0}}}\int_{0}^{1}\frac{d}{ds}\frac{y-sz}{|y-sz|^{2}}dsv_{0}\cdot\nabla\eta_{R_{0}}(z)dz$
$=-\frac{1}{2\pi}\int_{B_{2R_{0}}}\int_{0}^{1}(-z)\cdot\nabla\frac{y-sz}{|y-sz|^{2}}dsv_{0}\cdot\nabla\eta_{R_{0}}(z)dz.$
Hence,
$|w(y)|\leq$ $\frac{1}{2\pi}\int_{B_{2R_{0}}}\int_{0}^{1}|z|\frac{1}{|y-sz|^{2}}ds|v_{0}\cdot\nabla\eta_{R_{0}}|dz$
$\leq\frac{9R_{0}}{\pi}\int_{B_{2R_{0}}}\frac{1}{|y|^{2}}|v_{0}\cdot\nabla\eta_{R_{0}}|dz\leq\frac{C(A,R_{0})}{|y|^{2}},$
where we have used that $|y-sz|\geq|y|-s|z|\geq|y|-|z|\geq\frac{1}{3}|y|$ if $|y|\geq 3R_{0}$ and $|z|\leq 2R_{0}$. This shows that
$|w(y)|\leq\frac{C(A,R_{0})}{1+|y|^{2}}.$
The proof for $|\nabla^{k}w(y)|\leq\frac{C(A,\beta,k,R_{0})}{1+|y|^{k+2}}$, $k\geq 1$, is similar to the above calculation by taking the derivatives into the kernel and using the property $\int_{\mathbb{R}^{2}}v_{0}\cdot\nabla\eta_{R_{0}}=0$, we omit details.
To prove (2.18), we first note due to Lemma 2.1 and the definition of $\eta_{R_{0}}$ that
$|\eta_{R_{0}}v_{0}|\leq\frac{C(A)}{1+|y|}\mathbbm{1}_{B_{R_{0}}^{c}},~|\nabla^{k}(\eta_{R_{0}}v_{0})|\leq\frac{C(A,\beta)}{1+|y|^{1+\beta}}\mathbbm{1}_{B_{R_{0}}^{c}},~\textrm{ for }k=1,2,$
where $\mathbbm{1}$ is the characteristic function and $B_{R_{0}}^{c}$ is the complement of $B_{R_{0}}$. It is then easy to see that for any $r>2$ we have
$\|\eta_{R_{0}}v_{0}\|_{W^{2,r}(\mathbb{R}^{2})}\to 0\textrm{ as }R_{0}\to+\infty.$
Using the integral formula for $w$ and the Calderon–Zygmund estimates, we have for any $r>2$ that
$\|w\|_{W^{2,r}(\mathbb{R}^{2})}\leq c(r)\|\eta_{R_{0}}v_{0}\|_{W^{2,r}(\mathbb{R}^{2})}.$
It also follows from Morrey’s inequality that for any $r>2$
$\|w\|_{C^{1,\alpha}(\mathbb{R}^{2})}\leq c(r)\|w\|_{W^{2,r}(\mathbb{R}^{2})},$
where $\alpha=1-\frac{2}{r}$. Hence, we have
$\|w\|_{C^{1,\alpha}(\mathbb{R}^{2})}\leq c(r)\|\eta_{R_{0}}v_{0}\|_{W^{2,r}(\mathbb{R}^{2})}\to 0\textrm{ as }R_{0}\to+\infty.$
Choose $r=3$ in the above, for any $\epsilon>0$, we can then choose $R_{0}=R_{0}(A,\beta,\epsilon)$ large enough such that
$\|w\|_{C^{1,\frac{1}{3}}(\mathbb{R}^{2})}\leq C\|\eta_{R_{0}}v_{0}\|_{W^{2,3}(\mathbb{R}^{2})}\leq\frac{\epsilon}{2}.$
This $R_{0}=R_{0}(A,\beta,\epsilon)$ can already be large enough to make
$\|\eta_{R_{0}}v_{0}\|_{C^{1}(\mathbb{R}^{2})}\leq\frac{\epsilon}{2}.$
Combining the above two inequalities, we obtain (2.18). We conclude the whole proof. ∎
With this new profile $v_{1}=\eta_{R_{0}}v_{0}+w$, we define the modified version of $v_{\textrm{re}}$ as
(2.19) $v_{2}=v-v_{1},$
where $v$ is the solution to (2.4). Then $v_{2}$ satisfies
(2.20) $\left\{\begin{aligned} &-\Delta v_{2}-\frac{1}{2}v_{2}-\frac{1}{2}y\cdot\nabla v_{2}+v_{1}\cdot\nabla v_{2}+v_{2}\cdot\nabla v_{1}+v_{2}\cdot\nabla v_{2}+\nabla q_{2}=-v_{1}\cdot\nabla v_{1}-F(v_{0},w),\\ &{\rm div\thinspace}v_{2}=0.\end{aligned}\right$.
where
(2.21) $F(v_{0},w)$ $=-\Delta v_{1}-\frac{1}{2}v_{1}-\frac{1}{2}y\cdot\nabla v_{1}$
$=-\Delta(\eta_{R_{0}}v_{0}+w)-\frac{1}{2}(\eta_{R_{0}}v_{0}+w)-\frac{1}{2}y\cdot\nabla(\eta_{R_{0}}v_{0}+w)$
$=-2\nabla\eta_{R_{0}}\nabla v_{0}-\Delta\eta_{R_{0}}v_{0}-\frac{1}{2}y\cdot\nabla\eta_{R_{0}}v_{0}-\Delta w-\frac{1}{2}w-\frac{1}{2}y\cdot\nabla w.$
To show Theorem 2.1, we first prove a slightly weaker version for $v_{2}$. It says that $v_{2}$ is is $L^{p}$ integrable and hence $v_{\text{re}}$ is also $L^{p}$ integrable for any $1<p<2$.
Theorem 2.2.
For any $1<p<2$, we can choose $R_{0}=R_{0}(A,\beta,p)$ in the definition of $v_{1}$ (2.14) large enough, such that $v_{2}=v-v_{1}$, the corresponding solution to (2.20) satisfies
$\int_{\mathbb{R}^{2}}|v_{2}|^{2}+|\nabla v_{2}|^{2}dy\leq C(A,\beta,p),$
and
$\|v_{2}\|_{L^{p}(\mathbb{R}^{2})}\leq C(A,\beta,p).$
Especially, choosing $p=\frac{3}{2}$ and fix a $R_{0}=R_{0}(A,\beta)$ hence the modified profile $v_{1}$, then there exists a constant $C$ depending only on $A$ and $\beta$ such that
$\int_{\mathbb{R}^{2}}|v_{2}|^{2}+|\nabla v_{2}|^{2}dy\leq C(A,\beta).$
Proof.
Step 1: Estimate of $\|\nabla v_{2}\|_{L^{2}}$. Multiplying (2.20) by $v_{2}$, then integrating both sides, it leads to
(2.22) $\int_{\mathbb{R}^{2}}|\nabla v_{2}|^{2}dy+\int_{\mathbb{R}^{2}}(v_{2}\cdot\nabla v_{1})\cdot v_{2}dy+\int_{\mathbb{R}^{2}}(v_{1}\cdot\nabla v_{1})\cdot v_{2}dy=-\int_{\mathbb{R}^{2}}F(v_{0},w)v_{2}.$
Multiplying (2.20) by $v_{1}$, then integrating both sides, it leads to
(2.23) $\int_{\mathbb{R}^{2}}\partial_{i}v_{2,j}\partial_{i}v_{1,j}dy+\int_{\mathbb{R}^{2}}-\frac{1}{2}v_{1}\cdot v_{2}-\frac{1}{2}(y\cdot\nabla v_{2})\cdot v_{1}dy$
$~+\int_{\mathbb{R}^{2}}(v_{1}\cdot\nabla v_{2})\cdot v_{1}dy+\int_{\mathbb{R}^{2}}(v_{2}\cdot\nabla v_{2})\cdot v_{1}dy=-\int_{\mathbb{R}^{2}}F(v_{0},w)v_{1}.$
Here and in the following $v_{1,j}$ means the $j$-th component of $v_{1}$, $v_{2,j}$ means the $j$-th component of $v_{2}$ for $j=1,2$. Multiplying the equation of $v_{1}$ (2.21) by $v_{2}$, then integrating both sides, it leads to
$\int_{\mathbb{R}^{2}}\partial_{i}v_{2,j}\partial_{i}v_{1,j}dy-\int_{\mathbb{R}^{2}}\frac{1}{2}v_{1}\cdot v_{2}+\frac{1}{2}y_{i}\partial_{i}v_{1,j}v_{2,j}dy$
$=$ $\int_{\mathbb{R}^{2}}\partial_{i}v_{2,j}\partial_{i}v_{1,j}dy+\int_{\mathbb{R}^{2}}-\frac{1}{2}v_{1}\cdot v_{2}+v_{1}\cdot v_{2}+\frac{1}{2}y_{i}\partial_{i}v_{2,j}v_{1,j}dy=\int_{\mathbb{R}^{2}}F(v_{0},w)\cdot v_{2},$
where integration by parts has been used. This is just
(2.24) $\int_{\mathbb{R}^{2}}\partial_{i}v_{2,j}\partial_{i}v_{1,j}dy+\int_{\mathbb{R}^{2}}\frac{1}{2}v_{1}\cdot v_{2}+\frac{1}{2}(y\cdot\nabla v_{2})\cdot v_{1}dy=\int_{\mathbb{R}^{2}}F(v_{0},w)\cdot v_{2}.$
We note that
$\int_{\mathbb{R}^{2}}(v_{2}\cdot\nabla v_{2})\cdot v_{1}=\int_{\mathbb{R}^{2}}v_{2,i}\partial_{i}v_{2,j}v_{1,j}dy=-\int_{\mathbb{R}^{2}}v_{2,i}v_{2,j}\partial_{i}v_{1,j}dy=-\int_{\mathbb{R}^{2}}(v_{2}\cdot\nabla v_{1})\cdot v_{2},$
and
$\int_{\mathbb{R}^{2}}(v_{1}\cdot\nabla v_{2})\cdot v_{1}=\int_{\mathbb{R}^{2}}{v}_{1,i}\partial_{i}v_{2,j}v_{1,j}dy=-\int_{\mathbb{R}^{2}}{v}_{1,i}v_{2,j}\partial_{i}v_{1,j}dy=-\int_{\mathbb{R}^{2}}(v_{1}\cdot\nabla v_{1})\cdot v_{2}.$
Hence, adding (2.22)-(2.24) together, we have
(2.25) $\int_{\mathbb{R}^{2}}|\nabla v_{2}|^{2}dy+2\int_{\mathbb{R}^{2}}\partial_{i}v_{2,j}\partial_{i}v_{1,j}dy=-\int_{\mathbb{R}^{2}}F(v_{0},w)\cdot v_{1}$
$=$ $\int_{\mathbb{R}^{2}}\left(2\nabla\eta_{R_{0}}\nabla v_{0}+\Delta\eta_{R_{0}}v_{0}+\frac{1}{2}y\cdot\nabla\eta_{R_{0}}v_{0}+\Delta w+\frac{1}{2}w+\frac{1}{2}y\cdot\nabla w\right)\cdot v_{1}.$
For the right-hand side of above, we estimate with the help of Lemma 2.2 as follows:
$\int_{\mathbb{R}^{2}}\left(2\nabla\eta_{R_{0}}\nabla v_{0}+\Delta\eta_{R_{0}}v_{0}+\frac{1}{2}y\cdot\nabla\eta_{R_{0}}v_{0}\right)\cdot v_{1}$
$\leq$ $\int_{\mathbb{R}^{2}}\frac{C(A,R_{0})}{1+|y|}\left|2\nabla\eta_{R_{0}}\nabla v_{0}+\Delta\eta_{R_{0}}v_{0}+\frac{1}{2}y\cdot\nabla\eta_{R_{0}}v_{0}\right|$
$\leq$ $\int_{B_{2R_{0}}}\frac{C(A,R_{0})}{1+|y|^{2}}=C(A,R_{0}),$
and it follows from (2.17) that
$\int_{\mathbb{R}^{2}}\left(\frac{1}{2}w+\frac{1}{2}y\cdot\nabla w\right)\cdot v_{1}$
$\leq$ $C(A,R_{0})\int_{\mathbb{R}^{2}}\frac{1}{1+|y|}\left|\frac{1}{2}w+\frac{1}{2}y\cdot\nabla w\right|$
$\leq$ $\int_{\mathbb{R}^{2}}\frac{C(A,\beta,R_{0})}{1+|y|^{3}}=C(A,\beta,R_{0}).$
Finally, using Lemma 2.2 again, we have
$\int_{\mathbb{R}^{2}}\Delta w\cdot v_{1}=-\int_{\mathbb{R}^{2}}\nabla w\cdot\nabla v_{1}\leq\int_{\mathbb{R}^{2}}\frac{C(A,\beta,R_{0})}{1+|y|^{4}}=C(A,\beta,R_{0}).$
Hence, go back to (2.25), we have
$\int_{\mathbb{R}^{2}}|\nabla v_{2}|^{2}dy+2\int_{\mathbb{R}^{2}}\partial_{i}v_{2,j}\partial_{i}v_{1,j}dy=-\int_{\mathbb{R}^{2}}F(v_{0},w)v_{1}\leq C(A,\beta,R_{0}),$
This together with Young’s inequality and (2.16) yields that
(2.26) $\int_{\mathbb{R}^{2}}|\nabla v_{2}|^{2}dy\leq 4\int_{\mathbb{R}^{2}}|\nabla v_{1}|^{2}dy-2\int_{\mathbb{R}^{2}}F(v_{0},w)v_{1}\leq C(A,\beta,R_{0}),$
where we have used
(2.27) $\int_{\mathbb{R}^{2}}|\nabla v_{1}|^{2}dy\leq\int_{\mathbb{R}^{2}}\frac{C(A,\beta,R_{0})}{1+|y|^{2+2\beta}}dy\leq C(A,\beta,R_{0}).$
This is our initial Dirichlet energy estimate.
Step 2: Estimate of $\|v_{2}\|_{L^{p}}$. We multiply the equation of $v_{2}$ (2.20) by $|v_{2}|^{p-2}v_{2}$, where $1<p<2$, and integrate on both sides, we have
(2.28) $\int_{\mathbb{R}^{2}}\left(-\Delta v_{2}-\frac{1}{2}v_{2}-\frac{1}{2}y\cdot\nabla v_{2}+v_{1}\cdot\nabla v_{2}+v_{2}\cdot\nabla v_{2}+v_{2}\cdot\nabla v_{1}+\nabla q_{2}\right)\cdot|v_{2}|^{p-2}v_{2}$
$\qquad\qquad=-\int_{\mathbb{R}^{2}}\left(v_{1}\cdot\nabla v_{1}-F(v_{0},w)\right)\cdot|v_{2}|^{p-2}v_{2}.$
We consider the left-hand side of (2.28) first.
Estimates for the first five terms on left-hand side of (2.28). For the first term, we have
$\int_{\mathbb{R}^{2}}-\Delta v_{2}\cdot|v_{2}|^{p-2}v_{2}=\int_{\mathbb{R}^{2}}-\Delta v_{2,i}|v_{2}|^{p-2}v_{2,i}$
$=$ $\int_{\mathbb{R}^{2}}\nabla v_{2,i}\cdot|v_{2}|^{p-2}\nabla v_{2,i}+\partial_{j}v_{2,i}(p-2)|v_{2}|^{p-3}\partial_{j}\sqrt{v_{2,1}^{2}+v_{2,2}^{2}}v_{2,i}$
$=$ $\int_{\mathbb{R}^{2}}|\nabla v_{2}|^{2}|v_{2}|^{p-2}+\partial_{j}v_{2,i}(p-2)|v_{2}|^{p-3}\frac{v_{2,1}\partial_{j}v_{2,1}+v_{2,2}\partial_{j}v_{2,2}}{\sqrt{v_{2,1}^{2}+v_{2,2}^{2}}}v_{2,i}$
$=$ $\int_{\mathbb{R}^{2}}|\nabla v_{2}|^{2}|v_{2}|^{p-2}+(p-2)|v_{2}|^{p-4}(v_{2,1}\partial_{j}v_{2,1}+v_{2,2}\partial_{j}v_{2,2})v_{2,i}\partial_{j}v_{2,i}$
$=$ $\int_{\mathbb{R}^{2}}|\nabla v_{2}|^{2}|v_{2}|^{p-2}+\sum_{j=1}^{2}(p-2)|v_{2}|^{p-4}(v_{2,i}\partial_{j}v_{2,i})^{2}$
$\geq$ $\int_{\mathbb{R}^{2}}|\nabla v_{2}|^{2}|v_{2}|^{p-2}+(p-2)|v_{2}|^{p-2}|\nabla v_{2}|^{2}=\int_{\mathbb{R}^{2}}(p-1)|\nabla v_{2}|^{2}|v_{2}|^{p-2},$
where in the last inequality, we have used that $p<2$. In above, $v_{2,i}$ is the $i$-th component of $v_{2}$ and $\partial_{j}v_{2,i}=\frac{\partial v_{2,i}}{\partial x_{j}}$ for $i,j=1,2$. Also, we have
$\int_{\mathbb{R}^{2}}-\frac{1}{2}y\cdot\nabla v_{2}\cdot|v_{2}|^{p-2}v_{2}$ $=\int_{\mathbb{R}^{2}}-\frac{1}{4}y\cdot\nabla|v_{2}|^{2}|v_{2}|^{p-2}$
$=\int_{\mathbb{R}^{2}}-\frac{1}{4}y\cdot\nabla|v_{2}|^{2}(|v_{2}|^{2})^{\frac{p-2}{2}}$
$=\int_{\mathbb{R}^{2}}-\frac{1}{4}y\cdot\nabla(|v_{2}|^{2})^{\frac{p}{2}}\frac{2}{p}$
$=\frac{1}{p}\int_{\mathbb{R}^{2}}|v_{2}|^{p}$
And using integration by parts, we have by using ${\rm div\thinspace}v_{1}=0$ that
[…]
Similarly, we obtain
[…]
From the above estimates, we have
(2.29) $\int_{\mathbb{R}^{2}}(-\Delta v_{2}-\frac{1}{2}v_{2}-\frac{1}{2}y\cdot\nabla v_{2}+v_{1}\cdot\nabla v_{2}+v_{2}\cdot\nabla v_{2})\cdot|v_{2}|^{p-2}v_{2}$
$\geq$ $\int_{\mathbb{R}^{2}}(p-1)|\nabla v_{2}|^{2}|v_{2}|^{p-2}+\left(\frac{1}{p}-\frac{1}{2}\right)|v_{2}|^{p}.$
Importantly, one note that p1>0p-1>0 and $\frac{1}{p}-\frac{1}{2}>0$ due to $1<p<2$.
Estimates for the last two terms on left-hand side of (2.28). For the second last term on the left-hand side of (2.28), we have
(2.30) $\left|\int_{\mathbb{R}^{2}}(v_{2}\cdot\nabla v_{1})\cdot|v_{2}|^{p-2}v_{2}\right|\leq\|\nabla v_{1}\|_{L^{\infty}}\int_{\mathbb{R}^{2}}|v_{2}|^{p}.$
To estimate the last term on the left-hand side of (2.28), we first note that
(2.31) $-\Delta q_{2}$ = $=\textrm{ div }(v_{1}\cdot\nabla v_{2}+v_{2}\cdot\nabla v_{1}+v_{2}\cdot\nabla v_{2}+v_{1}\cdot\nabla v_{1})-{\rm div\thinspace}F(v_{0},w)$
= $=\textrm{ div }{\rm div\thinspace}((v_{1}+v_{2})\otimes(v_{1}+v_{2})),$
since due to ${\rm div\thinspace}v_{1}=0$ and (2.21) that
${\rm div\thinspace}F(v_{0},w)={\rm div\thinspace}\left(-\Delta v_{1}-\frac{1}{2}v_{1}-\frac{1}{2}y\cdot\nabla v_{1}\right)=0.$
It follows from the Calderon–Zygmund theory that there exists a constant $c(p)$ such that
(2.32) $\|\nabla q_{2}\|_{L^{p}}$ $\leq c(p)\left(\|v_{1}\nabla v_{1}\|_{L^{p}}+\|v_{2}\nabla v_{1}\|_{L^{p}}\right$.
$\qquad\qquad+\left.\|v_{1}\nabla v_{2}\|_{L^{p}}+\|v_{2}\nabla v_{2}\|_{L^{p}}\right)$
We now fix some $p\in(1,2)$, let $\epsilon$ in Lemma 2.2 be small enough, say
$\epsilon\leq\min\left\{\frac{1}{10c(p)},\frac{1}{10}\right\}\cdot\left(\frac{1}{p}-\frac{1}{2}\right),$
by choosing
$R_{0}=R_{0}(A,\beta,\epsilon)=R_{0}(A,\beta,p)$
in (2.14) large enough. It follows that
$\|v_{1}\|_{L^{\infty}},~\|\nabla v_{1}\|_{L^{\infty}}\leq\epsilon\leq\min\left\{\frac{1}{10c(p)},\frac{1}{10}\right\}\cdot\left(\frac{1}{p}-\frac{1}{2}\right).$
Then from (2.26) we have
$\int_{\mathbb{R}^{2}}|\nabla v_{2}|^{2}dy\leq 4\int_{\mathbb{R}^{2}}|\nabla v_{1}|^{2}dy-2\int_{\mathbb{R}^{2}}F(v_{0},w)v_{1}\leq C(A,\beta,p).$
We can now estimate the last term on the left-hand side of (2.28) using (2.32) as follows:
(2.33) $\quad\int_{\mathbb{R}^{2}}(\nabla q_{2})\cdot|v_{2}|^{p-2}v_{2}\leq\|v_{2}\|_{L^{p}}^{p-1}\|\nabla q_{2}\|_{L^{p}}$
$\leq c(p)\|v_{2}\|_{L^{p}}^{p-1}\left(\|v_{1}\nabla v_{1}\|_{L^{p}}+\|v_{2}\nabla v_{1}\|_{L^{p}}+\|v_{1}\nabla v_{2}\|_{L^{p}}+\|v_{2}\nabla v_{2}\|_{L^{p}}\right)$
$\leq c(p)\|v_{2}\|_{L^{p}}^{p-1}\left(C(A,\beta,p)+\frac{1}{10c(p)}\left(\frac{1}{p}-\frac{1}{2}\right)\|v_{2}\|_{L^{p}}+\|v_{1}\|_{L^{\frac{2p}{2-p}}}\|\nabla v_{2}\|_{L^{2}}+\|\nabla v_{2}\|_{L^{2}}\|v_{2}\|_{L^{\frac{2p}{2-p}}}\right)$
$\leq c(p)\|v_{2}\|_{L^{p}}^{p-1}\left(C(A,\beta,p)+\frac{1}{10c(p)}\left(\frac{1}{p}-\frac{1}{2}\right)\|v_{2}\|_{L^{p}}+\|v_{1}\|_{L^{\frac{2p}{2-p}}}\|\nabla v_{2}\|_{L^{2}}+c_{1}(p)\|\nabla v_{2}\|_{L^{2}}^{1+\frac{p}{2}}\|v_{2}\|_{L^{p}}^{1-\frac{p}{2}}\right)$
$\leq c(p)\|v_{2}\|_{L^{p}}^{p-1}\left(C(A,\beta,p)+\frac{1}{10c(p)}\left(\frac{1}{p}-\frac{1}{2}\right)\|v_{2}\|_{L^{p}}+C(A,\beta,p)+C(A,\beta,p)\|v_{2}\|_{L^{p}}^{1-\frac{p}{2}}\right)$
$\leq C(A,\beta,p)+\frac{1}{5}\left(\frac{1}{p}-\frac{1}{2}\right)\|v_{2}\|_{L^{p}}^{p},$
where we have used Gagliardo–Nirenberg interpolation inequality
$\|v_{2}\|_{L^{\frac{2p}{2-p}}(\mathbb{R}^{2})}\leq c_{1}(p)\|\nabla v_{2}\|_{L^{2}(\mathbb{R}^{2})}^{\frac{p}{2}}\|v_{2}\|_{L^{p}(\mathbb{R}^{2})}^{1-\frac{p}{2}},~1<p<2,$
and also the fact that for any $1<p<2$,
$\|v_{1}\|_{L^{\frac{2p}{2-p}}}\leq C(A,\beta,p),~|v_{1}\nabla v_{1}|\leq\frac{C(A,\beta,R_{0})}{1+|y|^{2}}=\frac{C(A,\beta,p)}{1+|y|^{2}},~\|v_{1}\nabla v_{1}\|_{L^{p}}\leq C(A,\beta,p),$
which follows from Lemma 2.2. Equation (2.30) becomes
(2.34) $\left|\int_{\mathbb{R}^{2}}(v_{2}\cdot\nabla v_{1})\cdot|v_{2}|^{p-2}v_{2}\right|\leq\frac{1}{10}\left(\frac{1}{p}-\frac{1}{2}\right)\int_{\mathbb{R}^{2}}|v_{2}|^{p}.$
Estimates for the right-hand side of (2.28). We now turn to estimate the right-hand side of (2.28). We claim that for any $1<p<2$,
(2.35) $\|F(v_{0},w)\|_{L^{p}}\leq C(A,\beta,p).$
The verification of this claim is a direct consequence of Lemma 2.2, and we leave it to the end of this proof. We have by using (2.35) that
(2.36) $\quad\int_{\mathbb{R}^{2}}(v_{1}\cdot\nabla v_{1}-F(v_{0},w))\cdot|v_{2}|^{p-2}v_{2}$
$\leq\int_{\mathbb{R}^{2}}C|v_{1}\cdot\nabla v_{1}|^{p}+C|F(v_{0},w)|^{p}+\frac{1}{10}\left(\frac{1}{p}-\frac{1}{2}\right)|v_{2}|^{p}$
$\leq C(A,\beta,p)+\frac{1}{10}\left(\frac{1}{p}-\frac{1}{2}\right)\int_{\mathbb{R}^{2}}|v_{2}|^{p}.$
Combining (2.29), (2.33), (2.34) and (2.36), we have
$\int_{\mathbb{R}^{2}}|\nabla v_{2}|^{2}|v_{2}|^{p-2}+|v_{2}|^{p}\leq C(A,\beta,p).$
The application of the Gagliardo–Nirenberg interpolation inequality also gives that
$\|v_{2}\|_{L^{2}(\mathbb{R}^{2})}\leq\|v_{2}\|_{L^{p}(\mathbb{R}^{2})}^{\frac{p}{2}}\|\nabla v_{2}\|_{L^{2}(\mathbb{R}^{2})}^{1-\frac{p}{2}}\leq C(A,\beta,p).$
It remains to prove the claim of (2.35). Using (2.21), we have
$\int_{\mathbb{R}^{2}}|F(v_{0},w)|^{p}$
$\leq$ $C\int_{\mathbb{R}^{2}}\left|2\nabla\eta_{R_{0}}\nabla v_{0}+\Delta\eta_{R_{0}}v_{0}+\frac{1}{2}y\cdot\nabla\eta_{R_{0}}v_{0}+\frac{1}{2}w+\frac{1}{2}y\cdot\nabla w\right|^{p}$
$+\int_{\mathbb{R}^{2}}|\Delta w|^{p}\leq C(A,\beta,p),$
where we have used Lemma 2.2 to have for any $1<p<2$ that
$\int_{\mathbb{R}^{2}}|\Delta w|^{p}\\ \leq C(A,\beta,p),$
and that
$\left|2\nabla\eta_{R_{0}}\nabla v_{0}+\Delta\eta_{R_{0}}v_{0}+\frac{1}{2}y\cdot\nabla\eta_{R_{0}}v_{0}+\frac{1}{2}w+\frac{1}{2}y\cdot\nabla w\right|\leq\frac{C(A,\beta,p)}{1+|y|^{2}}.$
This finishes the proof. ∎
Remark 2.2.
Throughout the proof, the Dirichlet energy estimate of $v_{1}$ (or equivalently Dirichlet energy estimate of $v_{0}$) (2.27) is the only place where we use the Hölder continuity exponent $\beta>0$.
We next prove Theorem 2.1 using the $H^{2}$-estimate of $v_{\text{re}}$ to be proved in Proposition 3.2.
Proof of Theorem 2.1 .
As $u_{\text{re}}(\cdot,1)=v_{\text{re}}(\cdot)$, we just prove the statement for $v_{\text{re}}$. Using the standard bootstrapping argument, along with the regularity estimates of the Stokes equations, one can show that $v_{\text{re}}\in H^{1}(\mathbb{R}^{2})$ is indeed smooth. For example, one can refer to [28, Theorem 3.8] for a detailed presentation of the bootstrapping argument. This also implies the smoothness of $u(x,1)$ as $e^{\Delta}u_{0}$ is already smooth for $t>0$. Due to the definitions of $v_{\textrm{re}}$ (2.6) and $v_{2}$ (2.19), we have
$v_{\textrm{re}}-v_{2}=v_{1}-v_{0}=(\eta_{R_{0}}-1)v_{0}+w,$
Lemmas 2.1 and 2.2 together with Theorem 2.2 tells us that for $1<p<2$,
(2.37) $\|v_{\text{re}}\|_{L^{p}(\mathbb{R}^{2})}\leq C(A,\beta,p).$
and that
$\|v_{\text{re}}\|_{H^{1}(\mathbb{R}^{2})}\leq C(A,\beta).$
To show higher integrability, we can use estimates for elliptic equations to show that indeed $v_{\textrm{re}}$ belongs to $H^{2}(\mathbb{R}^{2})$, which will be done in Proposition 3.2. It follows from Proposition 3.2 that we have
(2.38) $\|v_{\text{re}}\|_{L^{\infty}(\mathbb{R}^{2})}\leq C(A,\beta).$
Now, (2.38) combines with (2.37) and the interpolation inequality implies that for ant $p>1$
$\|v_{\text{re}}\|_{L^{p}(\mathbb{R}^{2})}\leq C(A,\beta,p).$
Using the above and noting that
$\Delta q=\operatorname{div}(v_{\text{re}}\cdot\nabla v_{\text{re}}+v_{0}\cdot\nabla v_{\text{re}}+v_{\text{re}}\cdot\nabla v_{0}+v_{0}\cdot\nabla v_{0})=\textrm{ div }{\rm div\thinspace}((v_{0}+v_{\text{re}})\otimes(v_{0}+v_{\text{re}})),$
the Calderon–Zygmund estimates give that for any $p>1$ that
$\|q\|_{L^{p}(\mathbb{R}^{2})}\leq C(A,\beta,p).$
Similarly, by Calderon–Zygmund estimates, we have
$\|\nabla q\|_{L^{2}(\mathbb{R}^{2})}\leq C(A,\beta).$
3. Weighted Energy Estimates and Pointwise Decay
In this section, we establish the pointwise decay estimates for the remainder part $v_{\text{re}}$. For the three-dimensional case, Jia and Šverák [18] first utilized localized smoothing estimates for local Leray solutions to derive pointwise decay. Here, we adopt the methods developed in [22], and establish weighted energy estimates to $\sqrt{1+|y|^{2}}v_{\text{re}}$ in $H^{2}(\mathbb{R}^{2})$, which subsequently allows us to derive the desired pointwise decay via Sobolev embeddings. That is
$|v_{\text{re}}(y)|\leq\frac{C}{1+|y|}\textrm{ for all }y\in\mathbb{R}^{2}.$
With this decay estimate in hand, we can improve the estimates by using the linear theory of Stokes and the heat equation to have that
$|v_{\text{re}}(y)|\leq\frac{C}{1+|y|^{2}}\textrm{ for all }y\in\mathbb{R}^{2}.$
Furthermore, we will prove the optimal decay for $v_{\text{re}}$ if the initial data $u_{0}\in C^{2}_{loc}$.
In light of the global energy estimates established in Theorem 2.1, we do not need to work for $v_{2}$ and $v_{1}$ anymore, which are modified versions of $v_{\text{re}}$ and $v_{0}$. Hence, in this section, we directly work on $v_{\text{re}}$ and $v_{0}$. For convenience, we recall that the governing equation for the remainder $v_{\text{re}}$ is given by:
(3.1) $\left\{\begin{aligned} &-\Delta v_{\text{re}}-\frac{1}{2}v_{\text{re}}-\frac{1}{2}y\cdot\nabla v_{\text{re}}+v_{0}\cdot\nabla v_{\text{re}}+v_{\text{re}}\cdot\nabla v_{0}+v_{\text{re}}\cdot\nabla v_{\text{re}}+\nabla q=-v_{0}\cdot\nabla v_{0},\\ &{\rm div\thinspace}v_{\text{re}}=0.\end{aligned}\right$.
And we also recall that $v_{0}$ is defined in (2.5) and solves (2.7).
3.1. Weighted $H^{1}$-Estimates
As a first step towards the higher-order estimates, we establish a weighted $H^{1}$ bound for the remainder term $v_{\text{re}}$.
Proposition 3.1 (Weighted $H^{1}$ Estimate).
Let the divergence-free, self-similar initial data $u_{0}(x)\in C^{0,\beta}_{loc}$ be given, and that
$A=\|u_{0}\|_{C^{0,\beta}(S^{1})}<+\infty.$
Assume $u$ is a self-similar solution to (1.1) in the sense of Definition 1.1. Consider the corresponding solution $(v_{\text{re}},q)\in H^{1}(\mathbb{R}^{2})\times L^{2}(\mathbb{R}^{2})$ to (3.1), then we have
(3.2) $\|\sqrt{1+|y|^{2}}v_{\text{re}}\|_{H^{1}(\mathbb{R}^{2})}\leq C(A,\beta).$
Proof.
Throughout this proof, we use $C$ to denote a constant depending on $A$ and $\beta$ that may be different from line to line. We first note that
$|\nabla(\sqrt{1+|y|^{2}}v_{\text{re}})|\leq\frac{|y|}{\sqrt{1+|y|^{2}}}|v_{\text{re}}|+\sqrt{1+|y|^{2}}|\nabla v_{\text{re}}|,$
and hence
$|\nabla(\sqrt{1+|y|^{2}}v_{\text{re}})|^{2}\leq 2|v_{\text{re}}|^{2}+2(1+|y|^{2})|\nabla v_{\text{re}}|^{2}.$
As we already prove the $H^{1}$-estimates of $v_{\text{re}}$, it suffices to prove the $H^{1}$-estimates of $|y|v_{\text{re}}$.
For rigorous derivation, one should use a regularized weight $h_{\epsilon}(y)=\frac{|y|}{(1+\epsilon|y|^{2})^{1/2}}$ and test (2.20) with $\phi=|h_{\epsilon}|^{2}v_{\text{re}}$, then pass $\epsilon\to 0$. For clarity of the calculation, we present the estimate using the weight $h(y)=|y|$. We test (2.20) with $\phi=|y|^{2}v_{\text{re}}$.
Step 1: The Dissipation Term. We calculate the term $\int\Delta v_{\text{re}}\cdot(|y|^{2}v_{\text{re}})$. We have by direct calculations that
$-\int_{\mathbb{R}^{2}}\Delta v_{\text{re}}\cdot(|y|^{2}v_{\text{re}})\,dy$ $=\int_{\mathbb{R}^{2}}\nabla v_{\text{re}}:\nabla(|y|^{2}v_{\text{re}})\,dy$
$=\int_{\mathbb{R}^{2}}\nabla v_{\text{re}}:(|y|^{2}\nabla v_{\text{re}}+2y\otimes v_{\text{re}})\,dy$
$=\int_{\mathbb{R}^{2}}|y|^{2}|\nabla v_{\text{re}}|^{2}\,dy+\int_{\mathbb{R}^{2}}y\cdot\nabla(|v_{\text{re}}|^{2})\,dy$
$=\int_{\mathbb{R}^{2}}|y|^{2}|\nabla v_{\text{re}}|^{2}\,dy-\int_{\mathbb{R}^{2}}(\nabla\cdot y)|v_{\text{re}}|^{2}\,dy$
$=\||y|\nabla v_{\text{re}}\|_{L^{2}}^{2}-2\|v_{\text{re}}\|_{L^{2}}^{2}\geq\||y|\nabla v_{\text{re}}\|_{L^{2}}^{2}-C.$
Step 2: The Drift and Linear Damping Terms. The linear drift and damping part of the operator is $\mathcal{L}(v_{\text{re}})=-\frac{1}{2}v_{\text{re}}-\frac{1}{2}y\cdot\nabla v_{\text{re}}$. Testing this with $|y|^{2}v_{\text{re}}$, we compute:
(3.3) $I_{drift}=\int_{\mathbb{R}^{2}}\left(-\frac{1}{2}v_{\text{re}}-\frac{1}{2}y\cdot\nabla v_{\text{re}}\right)\cdot(|y|^{2}v_{\text{re}})\,dy.$
The first part is simply $\frac{1}{2}\int|y|^{2}|v_{\text{re}}|^{2}dy$. For the second part, we use integration by parts. Note that $\nabla v_{\text{re}}\cdot v_{\text{re}}=\frac{1}{2}\nabla(|v_{\text{re}}|^{2})$. Thus,
$-\frac{1}{2}\int_{\mathbb{R}^{2}}(y\cdot\nabla v_{\text{re}})\cdot(|y|^{2}v_{\text{re}})\,dy$ $=-\frac{1}{4}\int_{\mathbb{R}^{2}}y\cdot\nabla(|v_{\text{re}}|^{2})|y|^{2}\,dy$
$=\frac{1}{4}\int_{\mathbb{R}^{2}}|v_{\text{re}}|^{2}\nabla\cdot(y|y|^{2})\,dy.$
In 2D, we calculate the divergence explicitly:
$\nabla\cdot(y|y|^{2})=(\nabla\cdot y)|y|^{2}+y\cdot\nabla(|y|^{2})=2|y|^{2}+y\cdot(2y)=2|y|^{2}+2|y|^{2}=4|y|^{2}.$
Substituting this back, we have
$\frac{1}{2}\int_{\mathbb{R}^{2}}(y\cdot\nabla v_{\text{re}})\cdot(|y|^{2}v_{\text{re}})\,dy=\frac{1}{4}\int_{\mathbb{R}^{2}}|v_{\text{re}}|^{2}(4|y|^{2})\,dy=\int_{\mathbb{R}^{2}}|y|^{2}|v_{\text{re}}|^{2}\,dy.$
Combining with the damping term $\frac{1}{2}v_{\text{re}}$:
(3.4) $I_{drift}=-\frac{1}{2}\int_{\mathbb{R}^{2}}|y|^{2}|v_{\text{re}}|^{2}+\int_{\mathbb{R}^{2}}|y|^{2}|v_{\text{re}}|^{2}=\frac{1}{2}\int_{\mathbb{R}^{2}}|y|^{2}|v_{\text{re}}|^{2}\,dy.$
This positivity is specific to the special structure of the Leray system and ensures the weighted $L^{2}$ norm is controlled.
Step 3: Interaction Terms involving $v_{0}$. We estimate the term involving $v_{0}$. We have
$I_{int}$ $=\int_{\mathbb{R}^{2}}(v_{0}\cdot\nabla v_{\text{re}}+v_{\text{re}}\cdot\nabla v_{0})\cdot(|y|^{2}v_{\text{re}})dy$
$=-\int_{\mathbb{R}^{2}}v_{0}\cdot y|v_{\text{re}}|^{2}+v_{\text{re}}\cdot\nabla(|y|^{2}v_{\text{re}})\cdot v_{0}dy.$
By Lemma 2.1, $|v_{0}(y)|\leq C(1+|y|)^{-1}$. Hence,
[…]
$\leq C\int_{\mathbb{R}^{2}}(|v_{\text{re}}|^{2}+|\nabla v_{\text{re}}|^{2})dy+\frac{1}{10}\int_{\mathbb{R}^{2}}|y|^{2}|v_{\text{re}}|^{2}dy$
$\leq C+\frac{1}{10}\int_{\mathbb{R}^{2}}|y|^{2}|v_{\text{re}}|^{2}dy.$
Step 4: Nonlinear term $v_{\text{re}}\cdot\nabla v_{\text{re}}$ and pressure term. For the nonlinear term $v_{\text{re}}\cdot\nabla v_{\text{re}}$, we have
$\int_{\mathbb{R}^{2}}(v_{\text{re}}\cdot\nabla v_{\text{re}})\cdot|y|^{2}v_{\text{re}}$ $=\frac{1}{2}\int_{\mathbb{R}^{2}}|y|^{2}v_{\text{re}}\cdot\nabla|v_{\text{re}}|^{2}=-\int_{\mathbb{R}^{2}}(y\cdot v_{\text{re}})|v_{\text{re}}|^{2}$
$\leq C\int_{\mathbb{R}^{2}}|v_{\text{re}}|^{4}dy+\frac{1}{10}\int_{\mathbb{R}^{2}}|y|^{2}|v_{\text{re}}|^{2}dy$
$\leq C\|v_{\text{re}}\|_{H^{1}(\mathbb{R}^{2})}^{4}+\frac{1}{10}\int_{\mathbb{R}^{2}}|y|^{2}|v_{\text{re}}|^{2}dy$
$\leq C+\frac{1}{10}\int_{\mathbb{R}^{2}}|y|^{2}|v_{\text{re}}|^{2}dy.$
We have
$\|q\|_{L^{2}}\leq C\left(\|v_{0}\|_{L^{4}}+\|v_{\text{re}}\|_{L^{4}}\right)^{2}\leq C\left(\|v_{0}\|_{L^{4}}+\|v_{\text{re}}\|_{H^{1}}\right)^{2}\leq C.$
Hence, using Hölder inequality, we have
$\int_{\mathbb{R}^{2}}\nabla q\cdot|y|^{2}v_{\text{re}}$ $=-\int_{\mathbb{R}^{2}}2qv_{\text{re}}\cdot y\leq\frac{1}{10}\int_{\mathbb{R}^{2}}|y|^{2}|v_{\text{re}}|^{2}+C$
Step 5: The source terms. The source term decay fast enough to be integrable against $|y|^{2}$. Indeed, we have
$\quad\int_{\mathbb{R}^{2}}(-v_{0}\cdot\nabla v_{0})\cdot|y|^{2}v_{\text{re}}$ $\leq C\int_{\mathbb{R}^{2}}|\nabla v_{0}|^{2}+\frac{1}{10}\int_{\mathbb{R}^{2}}|y|^{2}|v_{\text{re}}|^{2}$
$\leq C+\frac{1}{10}\int_{\mathbb{R}^{2}}|y|^{2}|v_{\text{re}}|^{2},$
where we have used that $\int_{\mathbb{R}^{2}}|\nabla v_{0}|^{2}\leq C$.
Combining Steps 1-5, we obtain:
$\||y|\nabla v_{\text{re}}\|_{L^{2}}^{2}+\||y|v_{\text{re}}\|_{L^{2}}^{2}\leq C$
This proves the proposition. ∎
3.2. Weighted $H^{2}$ Estimates for $v_{\textnormal{re}}$
With the weighted $H^{1}$ bounds established in Proposition 3.1, we now seek higher-order regularity for the remainder $v_{\text{re}}$. In the two-dimensional setting, $H^{1}$ regularity is insufficient to ensure a pointwise estimate. And it is necessary to establish estimates in the weighted Sobolev space $H^{2}$. The main goal of this section is to prove the $H^{2}$-estimate for $\sqrt{1+|y|^{2}}v_{\text{re}}$, which implies that $\sqrt{1+|y|^{2}}v_{\text{re}}(y)$ is bounded. We begin with a lemma establishing the improved regularity of a general weak solution in the usual Sobolev space.
Lemma 3.1 ($H^{2}$-estimates).
Let $f(y)\in L^{2}(\mathbb{R}^{2})$ be given. Assume that $(V,P)$ is a weak solution to the following system:
$\begin{cases}-\Delta V+V\cdot\nabla V-\dfrac{1}{2}(y\cdot\nabla V+V)+\nabla P=f&\text{in }\mathbb{R}^{2},\\ \operatorname{div}V=0.\end{cases}$
Specifically, $V\in H^{1}(\mathbb{R}^{2})$, $|y|V(y)\in L^{2}(\mathbb{R}^{2})$, $P\in L^{2}(\mathbb{R}^{2})$, and for all vector fields $\varphi\in H^{1}(\mathbb{R}^{2})$ with $|y|\varphi(y)\in L^{2}(\mathbb{R}^{2})$, the following identity holds:
(3.45) $\int_{\mathbb{R}^{2}}\nabla V:\nabla\varphi\,dy-\frac{1}{2}\int_{\mathbb{R}^{2}}(y\cdot\nabla V+V)\cdot\varphi\,dy$
$=\int_{\mathbb{R}^{2}}P\operatorname{div}\varphi\,dy+\int_{\mathbb{R}^{2}}V\cdot\nabla\varphi\cdot V\,dy+\int_{\mathbb{R}^{2}}f\cdot\varphi\,dy.$
Then, there exists a universal positive constant $C$ such that
$\|V\|_{H^{2}(\mathbb{R}^{2})}\leq C\left(\|V\|_{H^{1}(\mathbb{R}^{2})}+\|\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{2}\|V\|_{L^{2}(\mathbb{R}^{2})}^{\frac{1}{2}}+\|f\|_{L^{2}(\mathbb{R}^{2})}\right).$
Proof.
Let $D_{k}^{h}u$ denote the difference quotient of $u$:
$D_{k}^{h}u(y)=\frac{u(y+h\mathbf{e}_{k})-u(y)}{h},\quad h\in\mathbb{R}\setminus\{0\}.$
Choosing $\varphi(y)=-D_{k}^{-h}(D_{k}^{h}V)$ as a test function in (3.45), we obtain
$-\int_{\mathbb{R}^{2}}\nabla V:\nabla D_{k}^{-h}(D_{k}^{h}V)\,dy$ $=-\int_{\mathbb{R}^{2}}\nabla V:D_{k}^{-h}(D_{k}^{h}\nabla V)\,dy$
$=\int_{\mathbb{R}^{2}}|D_{k}^{h}\nabla V|^{2}\,dy=\|D_{k}^{h}\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{2}.$
Integration by parts yields
$\quad\frac{1}{2}\int_{\mathbb{R}^{2}}(y\cdot\nabla V+V)\cdot D_{k}^{-h}(D_{k}^{h}V)\,dy$
$=-\frac{1}{2}\int_{\mathbb{R}^{2}}D_{k}^{h}(y\cdot\nabla V+V)\cdot D_{k}^{h}V\,dy$
$=-\frac{1}{2}\int_{\mathbb{R}^{2}}\left[(y+h\mathbf{e}_{k})\cdot\nabla D_{k}^{h}V+(D_{k}^{h}y)\cdot\nabla V+D_{k}^{h}V\right]\cdot D_{k}^{h}V\,dy$
$=-\frac{1}{2}\int_{\mathbb{R}^{2}}y\cdot\nabla D_{k}^{h}V\cdot D_{k}^{h}V\,dy-\frac{1}{2}\int_{\mathbb{R}^{2}}\partial_{y_{k}}V\cdot D_{k}^{h}V\,dy-\frac{1}{2}\int_{\mathbb{R}^{2}}|D_{k}^{h}V|^{2}\,dy$
$=-\frac{1}{2}\int_{\mathbb{R}^{2}}\partial_{y_{k}}V\cdot D_{k}^{h}V\,dy,$
where we used $\int y\cdot\nabla\psi\cdot\psi\,dy=-\int\psi^{2}\,dy$ for $\psi=D_{k}^{h}V$. For the pressure term, integration by parts gives
$\int_{\mathbb{R}^{2}}P\operatorname{div}D_{k}^{-h}(D_{k}^{h}V)\,dy=\int_{\mathbb{R}^{2}}D_{k}^{h}P\operatorname{div}D_{k}^{h}V\,dy=0,$
owing to the divergence-free condition $\operatorname{div}V=0$. Similarly, for the nonlinear term, we have
$-\int_{\mathbb{R}^{2}}V\cdot\nabla D_{k}^{-h}(D_{k}^{h}V)\cdot V\,dy$ $=\int_{\mathbb{R}^{2}}V(y+h\mathbf{e}_{k})\cdot\nabla D_{k}^{h}V(y)\cdot D_{k}^{h}V(y)\,dy$
$\quad+\int_{\mathbb{R}^{2}}D_{k}^{h}V(y)\cdot\nabla D_{k}^{h}V(y)\cdot V(y)\,dy$
$=\int_{\mathbb{R}^{2}}D_{k}^{h}V(y)\cdot\nabla D_{k}^{h}V(y)\cdot V(y)\,dy.$
Substituting all above into (3.45) leads to
(3.5) $\|D_{k}^{h}\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{2}$ $=-\int_{\mathbb{R}^{2}}f(y)D_{k}^{-h}D_{k}^{h}V(y)\,dy+\frac{1}{2}\int_{\mathbb{R}^{2}}\partial_{y_{k}}V\cdot D_{k}^{h}V\,dy$
$\quad+\int_{\mathbb{R}^{2}}D_{k}^{h}V\cdot\nabla D_{k}^{h}V\cdot V\,dy.$
It remains to estimate the right-hand side of the above and pass to the limit as $h\to 0$.
First, note that
$\int_{\mathbb{R}^{2}}D_{k}^{-h}(D_{k}^{h}V)\cdot D_{k}^{-h}(D_{k}^{h}V)\,dy$ $=-\int_{\mathbb{R}^{2}}(D_{k}^{h}V)\cdot D_{k}^{h}D_{k}^{-h}(D_{k}^{h}V)\,dy$
$=\int_{\mathbb{R}^{2}}D_{k}^{h}(D_{k}^{h}V)\cdot D_{k}^{h}(D_{k}^{h}V)\,dy$
$\leq C\|\nabla D_{k}^{h}V\|_{L^{2}(\mathbb{R}^{2})}^{2}.$
This is just
$\|D_{k}^{-h}D_{k}^{h}V\|_{L^{2}}^{2}\leq C\|\nabla D_{k}^{h}V\|_{L^{2}(\mathbb{R}^{2})}^{2}.$
This allows us to estimate the first term on the right-hand side of (3.5) as
$\int_{\mathbb{R}^{2}}fD_{k}^{-h}D_{k}^{h}V\,dy\leq C\|f\|_{L^{2}}^{2}+\frac{1}{10}\|\nabla D_{k}^{h}V\|_{L^{2}}^{2}.$
By Hölder’s inequality, the second term on the right-hand side of (3.5) satisfies
$\frac{1}{2}\int_{\mathbb{R}^{2}}\partial_{y_{k}}V\cdot D_{k}^{h}V\,dy\leq\|\partial_{y_{k}}V\|_{L^{2}(\mathbb{R}^{2})}\|D_{k}^{h}V\|_{L^{2}(\mathbb{R}^{2})}\leq C\|\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{2}.$
Using the Gagliardo–Nirenberg interpolation inequality, the last term on the right-hand side of (3.5) is bounded by
$\int_{\mathbb{R}^{2}}D_{k}^{h}V(y)\cdot\nabla D_{k}^{h}V(y)\cdot V(y)\,dy$ $\leq\|V\|_{L^{6}(\mathbb{R}^{2})}\|D_{k}^{h}V\|_{L^{3}(\mathbb{R}^{2})}\|D_{k}^{h}\nabla V\|_{L^{2}(\mathbb{R}^{2})}$
$\leq C\|V\|_{L^{6}(\mathbb{R}^{2})}\|D_{k}^{h}V\|_{L^{2}(\mathbb{R}^{2})}^{2/3}\|D_{k}^{h}\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{4/3}$
$\leq C\|V\|_{L^{2}(\mathbb{R}^{2})}^{1/3}\|\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{2/3}\|D_{k}^{h}V\|_{L^{2}(\mathbb{R}^{2})}^{2/3}\|D_{k}^{h}\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{4/3}$
$\leq C\|V\|_{L^{2}(\mathbb{R}^{2})}^{1/3}\|\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{4/3}\|D_{k}^{h}\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{4/3}$
$\leq C\|\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{4}\|V\|_{L^{2}(\mathbb{R}^{2})}+\frac{1}{10}\|D_{k}^{h}\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{2}.$
Collecting all estimates above yields
$\|D_{k}^{h}\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{2}\leq C\|\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{2}+C\|\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{4}\|V\|_{L^{2}(\mathbb{R}^{2})}+C\|f\|_{L^{2}(\mathbb{R}^{2})}^{2}.$
Taking $h\to 0$ in the above inequality, we are led to
$\|\nabla^{2}V\|_{L^{2}(\mathbb{R}^{2})}^{2}\leq C\|\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{2}+C\|\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{4}\|V\|_{L^{2}(\mathbb{R}^{2})}+C\|f\|_{L^{2}(\mathbb{R}^{2})}^{2}.$
Hence
$\|V\|_{H^{2}(\mathbb{R}^{2})}$ $=\|V\|_{H^{1}(\mathbb{R}^{2})}+\|\nabla^{2}V\|_{L^{2}(\mathbb{R}^{2})}$
$\leq C\left(\|V\|_{H^{1}(\mathbb{R}^{2})}+\|\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{2}\|V\|_{L^{2}(\mathbb{R}^{2})}^{\frac{1}{2}}+\|f\|_{L^{2}(\mathbb{R}^{2})}\right).$
This completes the proof of the lemma. ∎
Proposition 3.2.
Under the same assumption of Proposition 3.1, we have
$\|v_{\text{re}}\|_{H^{2}(\mathbb{R}^{2})}\leq C(A,\beta),$
and
$\|v_{\text{re}}\|_{L^{\infty}(\mathbb{R}^{2})}\leq C(A,\beta).$
Proof.
By Proposition 3.1, we have that $|y|v_{\text{re}}\in H^{1}(\mathbb{R}^{2})$. We check that the force
$f=-(v_{0}\cdot\nabla v_{\text{re}}+v_{\text{re}}\cdot\nabla v_{0}+v_{0}\cdot\nabla v_{0})\in L^{2}(\mathbb{R}^{2}),~\|f\|_{\mathbb{R}^{2}}\leq C(A,\beta)$
due to $v_{0}$ is smooth bounded and satisfies the decay estimates in Lemma 2.1, and that $v_{re}\in H^{1}(\mathbb{R}^{2})$. Applying Lemma 3.1 with this $f$, we obtain that
$\|v_{\text{re}}(y)\|_{H^{2}(\mathbb{R}^{2})}\leq C(A,\beta).$
By the Sobolev embedding, we have
$\|v_{\text{re}}\|_{L^{\infty}(\mathbb{R}^{2})}\leq C(A,\beta).$
With these regularity estimates in hand, we are going to show the $H^{2}$-estimate for $\sqrt{1+|y|^{2}}v_{\text{re}}$, which implies that $\sqrt{1+|y|^{2}}v_{\text{re}}(y)$ is bounded.
Proposition 3.3 (Weighted $H^{2}$-estimates).
Under the same assumption of Proposition 3.1, we have
(3.6) $\|\sqrt{1+|y|^{2}}v_{\text{re}}\|_{H^{2}(\mathbb{R}^{2})}\leq C(A,\beta).$
Proof.
Having established the $H^{1}$-estimate of $\sqrt{1+|y|^{2}}v_{\text{re}}$, we now focus on the second-order derivative estimates. First, we observe the following expansion:
$\partial_{ij}(\sqrt{1+|y|^{2}}v_{\text{re}})=$ $\sqrt{1+|y|^{2}}\partial_{ij}v_{\text{re}}+\frac{y_{i}\partial_{j}v_{\text{re}}+y_{j}\partial_{i}v_{\text{re}}}{\sqrt{1+|y|^{2}}}\ +\left(\frac{\delta_{ij}}{\sqrt{1+|y|^{2}}}-\frac{y_{i}y_{j}}{(1+|y|^{2})^{3/2}}\right)v_{\text{re}}.$
To obtain $L^{2}$-estimates for $\partial_{ij}(\sqrt{1+|y|^{2}}v_{\text{re}})$, it is sufficient to bound $\sqrt{1+|y|^{2}}\partial_{ij}v_{\text{re}}$ in $L^{2}$ as all the other terms are bounded in $L^{2}$ due to Proposition 3.2. Furthermore, also in view of the $H^{2}$-estimate for $v_{\text{re}}$ provided in Proposition 3.2, it is equivalent to control the term $|y|\nabla^{2}v_{\text{re}}$ in $L^{2}(\mathbb{R}^{2})$.
To ensure a rigorous derivation, one should choose $\varphi(y)=-D_{k}^{-h}(h_{\epsilon}^{2}D_{k}^{h}v_{\text{re}})=-D_{k}^{-h}h_{\epsilon}^{2}D_{k}^{h}v_{\text{re}}$ as a test function in (3.1), where $h_{\epsilon}(y)=\frac{|y|}{(1+\epsilon|y|^{2})^{1/2}}$, and subsequently pass to the limit $\epsilon\to 0$. For the sake of clarity in the following presentation, we perform the estimates using the formal weight $h(y)=|y|$ and the test function $\varphi(y)=-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}})$.
Below, we use $C$ to denote a constant depending on $A$ and $\beta$, which may vary from line to line. Testing (3.1) with $\varphi(y)=-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}})$, we obtain
(3.7) $\int_{\mathbb{R}^{2}}\nabla v_{\text{re}}:\nabla(-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}}))\,dy-\frac{1}{2}\int_{\mathbb{R}^{2}}v_{\text{re}}\cdot(-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}}))\,dy$
$-\frac{1}{2}\int_{\mathbb{R}^{2}}y\cdot\nabla v_{\text{re}}\cdot(-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}}))\,dy-\int_{\mathbb{R}^{2}}q\operatorname{div}(-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}}))\,dy$
$=\int_{\mathbb{R}^{2}}v_{\text{re}}\cdot\nabla(-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}}))\cdot v_{\text{re}}\,dy-\int_{\mathbb{R}^{2}}v_{0}\cdot\nabla v_{\text{re}}\cdot(-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}}))\,dy$
$-\int_{\mathbb{R}^{2}}v_{\text{re}}\cdot\nabla v_{0}\cdot(-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}}))\,dy-\int_{\mathbb{R}^{2}}v_{0}\cdot\nabla v_{0}\cdot(-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}}))\,dy.$
We now calculate the above by utilizing the properties of difference quotients and the product rule for the gradient operator. Integration by parts leads to
(3.8) $\quad\int_{\mathbb{R}^{2}}\nabla v_{\text{re}}:\nabla(-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}}))dy$
$=\int_{\mathbb{R}^{2}}D_{k}^{h}\nabla v_{\text{re}}:\nabla(|y|^{2}D_{k}^{h}v_{\text{re}})dy$
$=\int_{\mathbb{R}^{2}}|y|^{2}|D_{k}^{h}\nabla v_{\text{re}}|^{2}dy+2\int_{\mathbb{R}^{2}}(D_{k}^{h}\nabla v_{\text{re}}):(y\otimes D_{k}^{h}v_{\text{re}})dy$
$=\||y|D_{k}^{h}\nabla v_{\text{re}}\|^{2}_{L^{2}(\mathbb{R}^{2})}+2\int_{\mathbb{R}^{2}}|y|D_{k}^{h}\nabla v_{\text{re}}:\left(\frac{y}{|y|}\otimes D_{k}^{h}v_{\text{re}}\right)dy$
$\geq\frac{9}{10}\||y|\nabla D_{k}^{h}v_{\text{re}}\|_{L^{2}}^{2}-C,$
where the second term in above is estimated via the Hölder inequality and the Young’s inequality as follows:
$-2\int_{\mathbb{R}^{2}}|y|D_{k}^{h}\nabla v_{\text{re}}:\left(\frac{y}{|y|}\otimes D_{k}^{h}v_{\text{re}}\right)dy$
$\leq$ $C\|\nabla v_{\text{re}}\|_{L^{2}(\mathbb{R}^{2})}\||y|D_{k}^{h}\nabla v_{\text{re}}\|_{L^{2}(\mathbb{R}^{2})}$
$\leq$ $C\|\nabla v_{\text{re}}\|_{L^{2}(\mathbb{R}^{2})}^{2}+\frac{1}{10}\||y|D_{k}^{h}\nabla v_{\text{re}}\|_{L^{2}(\mathbb{R}^{2})}^{2}$
$\leq$ $C+\frac{1}{10}\||y|\nabla D_{k}^{h}v_{\text{re}}\|_{L^{2}}^{2}.$
Utilizing the divergence-free condition $\mathrm{div}\,v_{\text{re}}=0$, the second and third terms on the left-hand side of (3.7) are estimated as:
$-\frac{1}{2}\int_{\mathbb{R}^{2}}y\cdot\nabla v_{\text{re}}\cdot(-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}}))\,dy-\frac{1}{2}\int_{\mathbb{R}^{2}}v_{\text{re}}\cdot(-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}}))\,dy$
$=$ $-\frac{1}{2}\int_{\mathbb{R}^{2}}(y+h\mathbf{e}_{k})\cdot\nabla D_{k}^{h}v_{\text{re}}|y|^{2}D_{k}^{h}v_{\text{re}}\,dy-\frac{1}{2}\int_{\mathbb{R}^{2}}\partial_{y_{k}}v_{\text{re}}\cdot(|y|^{2}D_{k}^{h}v_{\text{re}})\,dy$
$-\frac{1}{2}\int_{\mathbb{R}^{2}}|y|^{2}D_{k}^{h}v_{\text{re}}D_{k}^{h}v_{\text{re}}\,dy$
$=$ $\frac{1}{2}\int_{\mathbb{R}^{2}}|y|^{2}D_{k}^{h}v_{\text{re}}\cdot D_{k}^{h}v_{\text{re}}\,dy-\frac{h}{2}\int_{\mathbb{R}^{2}}\partial_{y_{k}}D_{k}^{h}v_{\text{re}}|y|^{2}D_{k}^{h}v_{\text{re}}\,dy$
$-\frac{1}{2}\int_{\mathbb{R}^{2}}\partial_{y_{k}}v_{\text{re}}\cdot(|y|^{2}D_{k}^{h}v_{\text{re}})\,dy$
$=$ $\frac{1}{2}\int_{\mathbb{R}^{2}}|y|^{2}D_{k}^{h}v_{\text{re}}\cdot D_{k}^{h}v_{\text{re}}\,dy+\frac{h}{2}\int_{\mathbb{R}^{2}}y_{k}D_{k}^{h}v_{\text{re}}\cdot D_{k}^{h}v_{\text{re}}\,dy$
$-\frac{1}{2}\int_{\mathbb{R}^{2}}\partial_{y_{k}}v_{\text{re}}\cdot(|y|^{2}D_{k}^{h}v_{\text{re}})\,dy$
$\geq$ $\frac{1}{4}\int_{\mathbb{R}^{2}}|y|^{2}D_{k}^{h}v_{\text{re}}\cdot D_{k}^{h}v_{\text{re}}\,dy+\frac{h}{2}\int_{\mathbb{R}^{2}}y_{k}D_{k}^{h}v_{\text{re}}\cdot D_{k}^{h}v_{\text{re}}\,dy$
$-\int_{\mathbb{R}^{2}}|y|^{2}|\partial_{y_{k}}v_{\text{re}}|^{2}\,dy$
$\geq$ $\frac{1}{4}\int_{\mathbb{R}^{2}}|y|^{2}|D_{k}^{h}v_{\text{re}}|^{2}\,dy+\frac{h}{2}\int_{\mathbb{R}^{2}}y_{k}|D_{k}^{h}v_{\text{re}}|^{2}\,dy-C.$
For the pressure term, a straightforward calculation yields:
$\int_{\mathbb{R}^{2}}q\,\mathrm{div}(-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}}))\,dy$ $=\int_{\mathbb{R}^{2}}D_{k}^{h}q\,\mathrm{div}(|y|^{2}D_{k}^{h}v_{\text{re}})\,dy$
$=\int_{\mathbb{R}^{2}}D_{k}^{h}q\nabla|y|^{2}\cdot D_{k}^{h}v_{\text{re}}\,dy$
$=2\int_{\mathbb{R}^{2}}D_{k}^{h}qy\cdot(D_{k}^{h}v_{\text{re}})\,dy.$
$\leq C\|\nabla q\|_{L^{2}}^{2}+\frac{1}{20}\|yD_{k}^{h}v_{\text{re}}\|_{L^{2}}^{2}.$
Regarding the convection terms on the right-hand side of (3.7), integration by parts and using Proposition 3.2 gives
$\int_{\mathbb{R}^{2}}v_{\text{re}}\cdot\nabla(-D_{k}^{-h}|y|^{2}D_{k}^{h}v_{\text{re}})\cdot v_{\text{re}}dy$
$=$ $\int_{\mathbb{R}^{2}}v_{\text{re}}(y+h\mathbf{e}_{k})\cdot\nabla(|y|^{2}D_{k}^{h}v_{\text{re}})\cdot D_{k}^{h}v_{\text{re}}dy+\int_{\mathbb{R}^{2}}D_{k}^{h}v_{\text{re}}\cdot\nabla(|y|^{2}D_{k}^{h}v_{\text{re}})\cdot v_{\text{re}}dy$
$=$ $\int_{\mathbb{R}^{2}}v_{\text{re}}(y+h\mathbf{e}_{k})\cdot 2y(D_{k}^{h}v_{\text{re}})\cdot D_{k}^{h}v_{\text{re}}dy+\int_{\mathbb{R}^{2}}|y|^{2}v_{\text{re}}(y+h\mathbf{e}_{k})\cdot\nabla(D_{k}^{h}v_{\text{re}})\cdot D_{k}^{h}v_{\text{re}}dy$
$+\int_{\mathbb{R}^{2}}D_{k}^{h}v_{\text{re}}\cdot 2y(D_{k}^{h}v_{\text{re}})\cdot v_{\text{re}}dy+\int_{\mathbb{R}^{2}}|y|^{2}D_{k}^{h}v_{\text{re}}\cdot\nabla(D_{k}^{h}v_{\text{re}})\cdot v_{\text{re}}dy$
$\leq$ $4\|v_{\text{re}}\|_{L^{\infty}}\||y|D_{k}^{h}v_{\text{re}}\|_{L^{2}}\|D_{k}^{h}v_{\text{re}}\|_{L^{2}}+2\|v_{\text{re}}\|_{L^{\infty}}\||y|D_{k}^{h}v_{\text{re}}\|_{L^{2}}\||y|\nabla D_{k}^{h}v_{\text{re}}\|_{L^{2}}$
$\leq$ $C\||y|D_{k}^{h}v_{\text{re}}\|_{L^{2}}^{2}+\frac{1}{20}\||y|\nabla D_{k}^{h}v_{\text{re}}\|_{L^{2}}^{2}.$
Similarly, for the interaction terms involving $v_{0}$, we have:
$-\int_{\mathbb{R}^{2}}v_{0}\cdot\nabla v_{\text{re}}\cdot(-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}}))$
$=$ $-\int_{\mathbb{R}^{2}}v_{0}(y+h\mathbf{e}_{k})\cdot D_{k}^{h}\nabla v_{\text{re}}\cdot(|y|^{2}D_{k}^{h}v_{\text{re}})-\int_{\mathbb{R}^{2}}D_{k}^{h}v_{0}\cdot\nabla v_{\text{re}}\cdot(|y|^{2}D_{k}^{h}v_{\text{re}})$
$\leq$ $\||y|v_{0}\|_{L^{\infty}}\|D_{k}^{h}v_{\text{re}}\|_{L^{2}}\||y|D_{k}^{h}\nabla v_{\text{re}}\|_{L^{2}}+\||y|^{2}\nabla v_{0}\|_{L^{\infty}}\|\nabla v_{\text{re}}\|_{L^{2}}^{2}$
$\leq$ $C\||y|v_{0}\|_{L^{\infty}}^{2}\|\nabla v_{\text{re}}\|_{L^{2}}^{2}+\||y|^{2}\nabla v_{0}\|_{L^{\infty}}\|\nabla v_{\text{re}}\|_{L^{2}}^{2}+\frac{1}{20}\||y|D_{k}^{h}\nabla v_{\text{re}}\|_{L^{2}}^{2}$
$\leq$ $C+\frac{1}{10}\||y|D_{k}^{h}\nabla v_{\text{re}}\|_{L^{2}}^{2}.$
For the last two terms on the right-hand side of (3.7), we have
$-\int_{\mathbb{R}^{2}}v_{\text{re}}\cdot\nabla v_{0}\cdot(-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}}))dy$
$=$ $-\int_{\mathbb{R}^{2}}v_{\text{re}}(y+he_{k})\cdot D_{k}^{h}\nabla v_{0}\cdot(|y|^{2}D_{k}^{h}v_{\text{re}})dy-\int_{\mathbb{R}^{2}}D_{k}^{h}v_{\text{re}}\cdot\nabla v_{0}\cdot(|y|^{2}D_{k}^{h}v_{\text{re}})dy$
$\leq$ $C\||y|^{2}\nabla^{2}v_{0}\|_{L^{\infty}}\|v_{\text{re}}\|_{L^{2}}\|\nabla v_{\text{re}}\|_{L^{2}}+C\||y|^{2}\nabla v_{0}\|_{L^{\infty}}\|\nabla v_{\text{re}}\|_{L^{2}}^{2}\leq C$
and
$-\int_{\mathbb{R}^{2}}v_{0}\cdot\nabla v_{0}\cdot(-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}}))dy$
$=$ $-\int_{\mathbb{R}^{2}}v_{0}(y+he_{k})\cdot D_{k}^{h}\nabla v_{0}\cdot(|y|^{2}D_{k}^{h}v_{\text{re}})dy-\int_{\mathbb{R}^{2}}D_{k}^{h}v_{0}\cdot\nabla v_{0}\cdot(|y|^{2}D_{k}^{h}v_{\text{re}})dy$
$\leq$ $\||y|v_{0}\|_{L^{\infty}}\|\nabla^{2}v_{0}\|_{L^{2}}\|y\|D_{k}^{h}v_{\text{re}}\|_{L^{2}}+\||y|^{2}\nabla v_{0}\|_{L^{\infty}}\|\nabla v_{0}\|_{L^{2}}\|\nabla v_{\text{re}}\|_{L^{2}}\leq C.$
Collecting all estimates and taking $h\to 0$ we obtain that
(3.9) $\||y|\nabla^{2}v_{\text{re}}\|^{2}_{L^{2}(\mathbb{R}^{2})}\leq C.$
This concludes the proof. ∎
3.3. Weighted $H^{2}$-Estimates for $\nabla v_{\textnormal{re}}$
We continue to establish the $H^{2}$-estimates for $\nabla v_{\textnormal{re}}$. The main goal of this section is to prove the following.
Proposition 3.4.
Under the same assumption of Proposition 3.1, we have
$\|v_{\text{re}}\|_{H^{3}(\mathbb{R}^{2})}\leq C(A,\beta),~\|\sqrt{1+|y|^{2}}\nabla v_{\text{re}}\|_{H^{2}(\mathbb{R}^{2})}\leq C(A,\beta).$
First, we give a lemma considering $H^{2}$-estimates for the following general linear system, which is quite similar to Lemma 3.1.
Lemma 3.2 ($H^{2}$-estimates).
Let $f(y)\in L^{2}(\mathbb{R}^{2})$ and the divergence-free vector field $\overline{V}\in H^{2}(\mathbb{R}^{2})$ be given. Assume that $(V,P)$ is a weak solution to the following system:
$\begin{cases}-\Delta V+\overline{V}\cdot\nabla V-\dfrac{1}{2}(y\cdot\nabla V+V)+\nabla P=f&\text{in }\mathbb{R}^{2},\\ \operatorname{div}V=0.\end{cases}$
Specifically, $V\in H^{1}(\mathbb{R}^{2})$, $|y|V(y)\in L^{2}(\mathbb{R}^{2})$, $P\in L^{2}(\mathbb{R}^{2})$, and for all vector fields $\varphi\in H^{1}(\mathbb{R}^{2})$, the following identity holds:
(3.10) $\int_{\mathbb{R}^{2}}\nabla V:\nabla\varphi\,dy-\frac{1}{2}\int_{\mathbb{R}^{2}}(y\cdot\nabla V+V)\cdot\varphi\,dy$
$=\int_{\mathbb{R}^{2}}P\operatorname{div}\varphi\,dy+\int_{\mathbb{R}^{2}}\overline{V}\cdot\nabla\varphi\cdot V\,dy+\int_{\mathbb{R}^{2}}f\cdot\varphi\,dy.$
Then, there exists a positive constant $C$ such that
$\|V\|_{H^{2}(\mathbb{R}^{2})}\leq C\left(\|\overline{V}\|_{H^{2}(\mathbb{R}^{2})},\|V\|_{H^{1}(\mathbb{R}^{2})},\|f\|_{L^{2}(\mathbb{R}^{2})}\right).$
Proof.
Similar to the proof of Lemma 3.1, we choose $\varphi=-D_{k}^{-h}D_{k}^{h}V$. All the other terms are identical to those in Lemma 3.1, we only need to deal with the following term. Using ${\rm div\thinspace}\overline{V}=0$, we have that
$-\int_{\mathbb{R}^{2}}\overline{V}\cdot\nabla D_{k}^{-h}(D_{k}^{h}V)\cdot V\,dy$ $=\int_{\mathbb{R}^{2}}\overline{V}(y+h\mathbf{e}_{k})\cdot\nabla D_{k}^{h}V(y)\cdot D_{k}^{h}V(y)\,dy$
$\quad+\int_{\mathbb{R}^{2}}D_{k}^{h}\overline{V}(y)\cdot\nabla D_{k}^{h}V(y)\cdot V(y)\,dy$
$=\int_{\mathbb{R}^{2}}D_{k}^{h}\overline{V}(y)\cdot\nabla D_{k}^{h}V(y)\cdot V(y)\,dy$
$\leq C\|V\|_{L^{6}(\mathbb{R}^{2})}\|D_{k}^{h}\overline{V}\|_{L^{3}(\mathbb{R}^{2})}\|D_{k}^{h}\nabla V\|_{L^{2}(\mathbb{R}^{2})}$
$\leq\frac{1}{10}\|D_{k}^{h}\nabla V\|_{L^{2}(\mathbb{R}^{2})}^{2}+C\|\overline{V}\|_{H^{2}(\mathbb{R}^{2})}^{2}\|V\|_{H^{1}(\mathbb{R}^{2})}^{2}.$
We finish the proof. ∎
We are now in a position to prove Proposition 3.4.
Proof of Proposition 3.4.
Let $\mathbb{P}=\textrm{Id}-\nabla(\Delta)^{-1}{\rm div\thinspace}$ be the Leray-Hopf projection onto the divergence-free vector fields. We know that $v_{\text{re}}$ solves
(3.11) $-\Delta v_{\text{re}}$ $=\frac{1}{2}y\cdot\nabla v_{\text{re}}+\frac{1}{2}v_{\text{re}}-\mathbb{P}(v_{\text{re}}\cdot\nabla v_{\text{re}}-v_{0}\cdot\nabla v_{\text{re}}-v_{\text{re}}\cdot\nabla v_{0}-v_{0}\cdot\nabla v_{0})$
$=f.$
Thanks to Lemma 2.1, Propositions 3.2 and 3.3, it is easy to see that this $f\in H^{1}(\mathbb{R}^{2})$. By the classical elliptic regularity theory and the fact that $v_{\textrm{re}}\in H^{2}(\mathbb{R}^{2})$, we immediately obtain
$\|v_{\textrm{re}}\|_{H^{3}(\mathbb{R}^{2})}\leq C.$
Hence, we also have that
$\|\nabla v_{\textrm{re}}\|_{L^{\infty}(\mathbb{R}^{2})}\leq C.$
Let $V_{k}=\sqrt{1+|y|^{2}}\partial_{y_{k}}v_{\text{re}}$ and $q_{k}=\sqrt{1+|y|^{2}}\partial_{y_{k}}q$, we can find that
(3.12) $-\Delta V_{k}+v_{\text{re}}\cdot\nabla V_{k}-\frac{1}{2}y\cdot\nabla V_{k}-\frac{1}{2}V_{k}+\nabla q_{k}$
$=-\sqrt{1+|y|^{2}}\partial_{y_{k}}(v_{\text{re}}\cdot\nabla v_{0}+v_{0}\cdot\nabla v_{\text{re}}+v_{0}\cdot\nabla v_{0})-V_{k}\cdot\nabla v_{\text{re}}$
$\quad+v_{\text{re}}\cdot\frac{y}{\sqrt{1+|y|^{2}}}\partial_{y_{k}}v_{\text{re}}-\frac{1}{2}V_{k}-\frac{|y|^{2}}{2\sqrt{1+|y|^{2}}}\partial_{y_{k}}v_{\text{re}}$
$\quad-2\frac{y}{\sqrt{1+|y|^{2}}}\cdot\nabla\partial_{y_{k}}v_{\text{re}}-\frac{2+|y|^{2}}{(1+|y|^{2})^{3/2}}\partial_{y_{k}}v_{\text{re}}+\frac{y}{\sqrt{1+|y|^{2}}}\partial_{y_{k}}q.$
We see that
$\sqrt{1+|y|^{2}}\partial_{y_{k}}(v_{\text{re}}\cdot\nabla v_{0}+v_{0}\cdot\nabla v_{\text{re}}+v_{0}\cdot\nabla v_{0})\in L^{2}(\mathbb{R}^{2}),$
and also
$\|V_{k}\|_{L^{2}(\mathbb{R}^{2})}\leq\|\sqrt{1+|y|^{2}}\nabla v_{\text{re}}\|_{L^{2}(\mathbb{R}^{2})},$
and
$\left\|\frac{y}{\sqrt{1+|y|^{2}}}\cdot\nabla\partial_{y_{k}}v_{\text{re}}\right\|_{L^{2}(\mathbb{R}^{2})}\leq C\|v_{\text{re}}\|_{H^{2}(\mathbb{R}^{2})}.$
The other terms on the right-hand side of (3.12) can be seen belong to $L^{2}(\mathbb{R}^{2})$. Applying Lemma 3.2, we know that
$\|\sqrt{1+|y|^{2}}\nabla v_{\textrm{re}}\|_{H^{2}(\mathbb{R}^{2})}\leq C.$
3.4. Pointwise Decay estimates for the self-similar solution
The main result of this section is Theorem 3.2, which asserts the optimal decay rates for the remainder term $v_{\textrm{re}}$. Propositions 3.3 and 3.4 implies the following obvious poinitwise estimate, which will be improved through a kind a bootstrap argument using the estimates for the Stokes system.
Theorem 3.1 (Pointwise decay estimates).
Under the same assumption of Proposition 3.1, we have for all $y\in\mathbb{R}^{2}$ that
$|\nabla^{k}v_{\text{re}}(y)|\leq\frac{C(A,\beta)}{1+|y|},~k=0,1.$
Proof.
From Proposition 3.3, we have $\sqrt{1+|y|^{2}}v_{\text{re}}\in H^{2}(\mathbb{R}^{2})$. By the Sobolev embedding $H^{2}(\mathbb{R}^{2})\hookrightarrow L^{\infty}(\mathbb{R}^{2})$, we have:
$\|\sqrt{1+|y|^{2}}v_{\text{re}}\|_{L^{\infty}}\leq C\|\sqrt{1+|y|^{2}}v_{\text{re}}\|_{H^{2}}\leq C.$
This implies
$|v_{\text{re}}(y)|\leq\frac{C}{1+|y|}.$
Similarly, it follows from Proposition 3.4 and Sobolev embedding that
$|\nabla v_{\text{re}}(y)|\leq\frac{C}{1+|y|}.$
We are going to improve the decay rate in Theorem 3.1 using the estimates for the Stokes system. We first study a nonhomogeneous Stokes system with a singular force.
(3.13) $\left\{\begin{aligned} \partial_{t}\vartheta-\Delta\vartheta+\nabla\pi&=t^{-1}{\rm div\thinspace}_{x}F\left(\frac{x}{\sqrt{t}}\right),\\ {\rm div\thinspace}\vartheta&=0,\end{aligned}\right.\text{ in }\mathbb{R}^{2}\times(0,+\infty).$
We have the following lemma, which plays an important role in improving decay estimates for weak solutions of equation (3.1). This kind of lemma already appeared in [18, Lemma 4.1] in the 3D case; see also [28, Lemma 3.15]. We present the lemma for the 2D case; the proof is essentially the same.
Lemma 3.3.
If for all $x\in\mathbb{R}^{2}$, $F(x)$ satisfies $|F(x)|\leq\frac{C}{1+|x|^{2}}$, then the solution $\vartheta(x,t)$ to (3.13) is given by
(3.14) $\vartheta(x,t)=\int_{0}^{t}e^{(t-s)\Delta}\mathbb{P}{\rm div\thinspace}_{x}s^{-1}F\left(\frac{\cdot}{\sqrt{s}}\right)ds.$
Let $\Theta(x)=\vartheta(x,1)$, then
(3.15) $|\Theta(x)|\leq\frac{C}{1+|x|^{2}}.$
Moreover, if $F\in C^{0,1}(\mathbb{R}^{2})$ satisfies ${\rm div\thinspace}_{x}F(\frac{x}{\sqrt{t}})=t^{-1/2}f(\frac{x}{\sqrt{t}})$ with $|f(x)|\leq\frac{C}{1+|x|^{2}}$, we have
(3.16) $|\nabla\Theta(x)|\leq\frac{C}{1+|x|^{2}}.$
Proof.
It is easy to check $\vartheta$ can be represented as (3.14). The kernel of $e^{t\Delta}\mathbb{P}$, usually called the Oseen kernel, which we denote by $S(x,t)$ is known to have the following estimates ([9]):
$|\partial_{t}^{l}\nabla^{k}_{x}S(x,t)|\leq C(k,l)t^{-l}(|x|+\sqrt{t})^{-2-k},~k,l\in\mathbb{N}^{+}.$
Note that
$\Theta(x)=\vartheta(x,1)$ $=\int_{0}^{1}e^{(1-s)\Delta}\mathbb{P}{\rm div\thinspace}_{x}s^{-1}F\left(\frac{\cdot}{\sqrt{s}}\right)ds$
$=-\int_{0}^{1}\int_{\mathbb{R}^{2}}\nabla S(x-y,1-s)s^{-1}F\left(\frac{y}{\sqrt{s}}\right)dyds.$
Hence, based on the pointwise estimate of the Oseen kernel, we have
(3.17) $|\Theta(x)|\leq C\int_{0}^{1}\int_{\mathbb{R}^{2}}\frac{1}{(\sqrt{1-s}+|x-y|)^{3}}\frac{1}{(\sqrt{s}+|y|)^{2}}dyds.$
If $F\in C^{1}(\mathbb{R}^{2})$ satisfies ${\rm div\thinspace}_{x}F(\frac{x}{\sqrt{t}})=t^{-1/2}f(\frac{x}{\sqrt{t}})$ with $|f(x)|\leq C(1+|x|)^{-2}$, we have
$\Theta(x)=\vartheta(x,1)$ $=\int_{0}^{1}e^{(1-s)\Delta}\mathbb{P}s^{-\frac{3}{2}}f\left(\frac{\cdot}{\sqrt{s}}\right)ds$
$=\int_{0}^{1}\int_{\mathbb{R}^{2}}S(x-y,1-s)s^{-\frac{3}{2}}f\left(\frac{y}{\sqrt{s}}\right)dyds.$
Then
$\nabla\Theta(x)=\int_{0}^{1}\int_{\mathbb{R}^{2}}\nabla S(x-y,1-s)s^{-\frac{3}{2}}f\left(\frac{y}{\sqrt{s}}\right)dyds.$
Therefore, we obtain, based on the pointwise estimate of the Oseen kernel, that
(3.18) $|\nabla\Theta(x)|$ $\leq\int_{0}^{1}s^{-\frac{1}{2}}\int_{\mathbb{R}^{2}}\frac{1}{(\sqrt{1-s}+|x-y|)^{3}}\frac{1}{(\sqrt{s}+|y|)^{2}}dyds$
We now prove that
$|\nabla\Theta(x)|\leq\frac{C}{1+|x|^{2}}$
by using (3.18). The proof of $|\Theta(x)|\leq\frac{C}{1+|x|^{2}}$ using (3.17) is similar and indeed simpler and thus omitted. We have
$|\nabla\Theta(x)|$ $\leq\int_{0}^{1}s^{-\frac{1}{2}}\int_{\mathbb{R}^{2}}\frac{1}{(\sqrt{1-s}+|x-y|)^{3}}\frac{1}{(\sqrt{s}+|y|)^{2}}dyds$
$=\int_{0}^{1}s^{-\frac{1}{2}}\int_{|y|\leq\frac{|x|}{2}}+\int_{\frac{|x|}{2}\leq|y|\leq 2|x|}+\int_{|y|\geq 2|x|}\frac{1}{(\sqrt{1-s}+|x-y|)^{3}}\frac{1}{(\sqrt{s}+|y|)^{2}}dyds$
$=J_{1}+J_{2}+J_{3}.$
We only consider $|x|\geq 10$, the boundedness when $|x|\leq 10$ is clear. Then
$J_{1}\leq\int_{0}^{1}s^{-\frac{1}{2}}\int_{|y|\leq\frac{|x|}{2}}\frac{8}{|x|^{3}}\frac{1}{(\sqrt{s}+|y|)^{2}}dyds\leq C|x|^{-3}\ln(1+|x|).$
$J_{2}$ $\leq\int_{0}^{1}s^{-\frac{1}{2}}\int_{\frac{|x|}{2}\leq|y|\leq 2|x|}\frac{1}{(\sqrt{1-s}+|x-y|)^{3}}\frac{1}{(\sqrt{s}+|y|)^{2}}dyds$
$\leq C\frac{1}{|x|^{2}}\int_{0}^{1}s^{-\frac{1}{2}}\int_{\frac{|x|}{2}\leq|y|\leq 2|x|}\frac{1}{(\sqrt{1-s}+|x-y|)^{3}}dyds$
$\leq C\frac{1}{|x|^{2}}\int_{0}^{1}s^{-\frac{1}{2}}\int_{|z|\leq 3|x|}\frac{1}{(\sqrt{1-s}+|z|)^{3}}dzds$
$\leq C\frac{1}{|x|^{2}}\int_{0}^{1}s^{-\frac{1}{2}}(1-s)^{-\frac{1}{2}}ds\leq C\frac{1}{|x|^{2}}.$
Similarly, one can prove
$J_{3}$ $\leq\int_{0}^{1}s^{-\frac{1}{2}}\int_{|y|\geq 2|x|}\frac{1}{(\sqrt{1-s}+|x-y|)^{3}}\frac{1}{(\sqrt{s}+|y|)^{2}}dyds$
$\leq C\frac{1}{|x|^{2}}\int_{0}^{1}s^{-\frac{1}{2}}\int_{|y|\geq 2|x|}\frac{1}{(\sqrt{1-s}+|y|)^{3}}dyds$
$\leq C\frac{1}{|x|^{2}}\int_{0}^{1}s^{-\frac{1}{2}}(1-s)^{-\frac{1}{2}}\leq C\frac{1}{|x|^{2}}.$
We finish the proof. ∎
We now give the main estimates for this section.
Theorem 3.2 (Improved pointwise decay estimates).
Under the same assumption of Proposition 3.1, we have for all $x\in\mathbb{R}^{2}$ that
$|v_{\text{re}}(x)|\leq C(A,\beta)(1+|x|)^{-2}$
and that
$|\nabla v_{\text{re}}(x)|\leq C(A,\beta)(1+|x|)^{-(2+\beta)}.$
Moreover, if the initial data $u_{0}\in C^{2}(S^{1})$, we have the following optimal decay rate estimates
$|v_{\text{re}}(x)|\leq C(\|u_{0}\|_{C^{2}(S^{1})})(1+|x|)^{-3}\ln(1+|x|).$
Proof.
Let $\vartheta(x,t)$ = $\frac{1}{\sqrt{t}}v_{\textrm{re}}\left(\frac{x}{\sqrt{t}}\right)$. Then this $\vartheta(x,t)$ solves
(3.19) $\left\{\begin{aligned} \partial_{t}\vartheta-\Delta\vartheta+\nabla\pi&=t^{-1}{\rm div\thinspace}_{x}F\left(\frac{x}{\sqrt{t}}\right),\\ {\rm div\thinspace}\vartheta&=0,\end{aligned}\right.\text{ in }\mathbb{R}^{2}\times(0,+\infty).$
with
$F=(v_{\textrm{re}}+v_{0})\otimes(v_{\textrm{re}}+v_{0}).$
It follows from Lemma 2.1 and Theorem 3.1 that
$|F|\leq\frac{C(A,\beta)}{1+|x|^{2}}.$
Using Lemma 3.3, we have
$|\vartheta(x,1)|=|v_{\textrm{re}}(x)|\leq\frac{C(A,\beta)}{1+|x|^{2}}.$
Moreover,
$t^{-1}{\rm div\thinspace}_{x}F\left(\frac{x}{\sqrt{t}}\right)=t^{-\frac{3}{2}}f\left(\frac{x}{\sqrt{t}}\right),$
where
$f=(v_{\textrm{re}}+v_{0})\cdot\nabla(v_{\textrm{re}}+v_{0}).$
It also follows from Lemma 2.1 and Theorem 3.1 that
$|f(x,t)|\leq\frac{C(A,\beta)}{1+|x|^{2}}.$
By then, Lemma 3.3 gives
$|\nabla v_{\textrm{re}}(x)|\leq\frac{C(A,\beta)}{1+|x|^{2}}.$
It follows from Lemma 2.1 and the above that
$|f(x,t)|\leq\frac{C(A,\beta)}{1+|x|^{2+\beta}}.$
We then can improve the decay of $\nabla v_{\textrm{re}}$ as
$|\nabla v_{\textrm{re}}(x)|$ $\leq\left|\int_{0}^{1}\int_{\mathbb{R}^{2}}\nabla S(x-y,1-s)s^{-\frac{3}{2}}f\left(\frac{y}{\sqrt{s}}\right)dyds\right|$
$\leq C\int_{0}^{1}\int_{\mathbb{R}^{2}}\frac{1}{(\sqrt{1-s}+|x-y|)^{3}}s^{\frac{\beta-1}{2}}\frac{1}{(\sqrt{s}+|y|)^{2+\beta}}dyds$
$\leq\frac{C}{1+|x|^{2+\beta}},~0<\beta\leq 1.$
If the initial data has higher regularity with $\|u_{0}\|_{C^{2}(S^{1})}<+\infty$, then we can apply [8, Lemma 4] to the vector field $F=v_{0}\otimes v_{0}=e^{\Delta}u_{0}\otimes e^{\Delta}u_{0}$, to conclude that
$\vartheta_{1}(x,t)=\int_{0}^{t}e^{(t-s)\Delta}\mathbb{P}{\rm div\thinspace}_{x}s^{-1}(e^{\Delta}u_{0}\otimes e^{\Delta}u_{0})\left(\frac{\cdot}{\sqrt{s}}\right)ds$
at time $t=1$ is bounded by $C(|u_{0}\|_{C^{2}(S^{1})})(1+|x|)^{-3}$. The proof makes essential use of the cancellation properties of the Oseen kernel. On the other hand, all the other terms in $(v_{\textrm{re}}+v_{0})\otimes(v_{\textrm{re}}+v_{0})$ except $v_{0}\otimes v_{0}$ has bound $C(1+|x|)^{-3}$ at least. And their contribution in (3.19) can be bound by
$C\int_{0}^{1}\int_{\mathbb{R}^{2}}\frac{1}{(\sqrt{1-s}+|x-y|)^{3}}\frac{1}{(\sqrt{s}+|y|)^{3}}dyds\leq\frac{C}{1+|x|^{3}}.$
4. Proof of the main theorem
The results in Sections 2 and 3 implies the following a priori estimates for the self-similar solutions on $\mathbb{R}^{2}$.
Theorem 4.1.
Let divergence-free initial data $u_{0}$ on $\mathbb{R}^{2}$ be self-similar, and that $\int_{\partial B_{R}}u_{0}\cdot nd\sigma=0$,
$A=\|u_{0}\|_{C^{0,\beta}(S^{1})}<+\infty,$
Assume $u(x,t)$ be a self-similar solution to (1.1) in the sense of Definition 1.1. Then $v(x)=u(x,1)$ the solution profile at time t = 1, belongs to $C^{\infty}(\mathbb{R}^{2})$. Let $v_{0}=e^{\Delta}u_{0}$ and $v_{\textrm{re}}(x)=v(x)-v_{0}(x)$, then we have
$\left|v_{\textrm{re}}(x)\right|=\left|v(x)-v_{0}(x)\right|\leq\frac{C(A,\beta)}{1+|x|^{2}}.$
Also,
$|\nabla v_{\text{re}}(x)|\leq\frac{C(A,\beta)}{(1+|x|)^{2+\beta}}.$
We can now prove Theorem 1.1.
Proof.
We first set up the framework for the application of the Leray-Schauder fixed point theorem. Define the function space
(4.1) $X=\{{\boldsymbol{v}}:\ {\rm div\thinspace}{\boldsymbol{v}}=0,\ {\boldsymbol{v}}\in C_{loc}(\mathbb{R}^{2})\},$
equipped with the norm
$\|{\boldsymbol{v}}\|_{X}=\|\sqrt{1+|x|^{2}}{\boldsymbol{v}}(x)\|_{C(\mathbb{R}^{2})}.$
The function space $X$ is a Banach space. For each ${\boldsymbol{v}}\in X$, we define its self-similar extension by
$E{\boldsymbol{v}}(x,t)=t^{-1/2}{\boldsymbol{v}}(t^{-1/2}x),\quad(x,t)\in\mathbb{R}^{3}\times(0,\infty).$
We now define an operator $K:X\times[0,1]\to X$ by solving the nonhomogeneous Stokes system and restrict the solution at time $t=1$:
(4.2) $K({\boldsymbol{v}},\sigma):=-\Phi[(\sigma e^{t\Delta}u_{0}+E{\boldsymbol{v}})\otimes(\sigma e^{t\Delta}u_{0}+E{\boldsymbol{v}})]\big|_{t=1}.$
In the above, $\Phi$ is the solution operator of the nonhomogeneous Stokes system defined by formula (3.14), in which one takes
$\left.F=(\sigma e^{t\Delta}u_{0}+E{\boldsymbol{v}})\otimes(\sigma e^{t\Delta}u_{0}+E{\boldsymbol{v}})\right|_{t=1}.$
Note that for ${\boldsymbol{v}}\in X$ with $\|{\boldsymbol{v}}\|_{X}<M$ and $0\leq\sigma\leq 1$, the force $\left.F=(\sigma v_{0}+E{\boldsymbol{v}})\otimes(\sigma v_{0}+E{\boldsymbol{v}})\right|_{t=1}$ satisfies
$|F(x)|\leq\frac{C(M,\|v_{0}\|_{X})}{1+|x|^{2}}\leq\frac{C(A,M)}{1+|x|^{2}}.$
It follows from Lemma 3.3 that
$|K({\boldsymbol{v}},\sigma)(x,1)|\leq\frac{C(A,M)}{1+|x|^{2}}\leq C(A,M)$
hence, $K({\boldsymbol{v}},\sigma)\in X$ and $\|K({\boldsymbol{v}},\sigma)\|_{X}\leq C(A,M)$. Thus $K$ indeed maps bounded sets in $X\times[0,1]$ into bounded sets in $X$.
Furthermore, $K$ is compact because its main term $\Phi[(\sigma e^{t\Delta}u_{0})\otimes(\sigma e^{t\Delta}u_{0})]\big|_{t=1}=\sigma^{2}\Phi[(e^{t\Delta}u_{0})\otimes(e^{t\Delta}u_{0})]\big|_{t=1}$ has a one-dimensional range. While the other terms of $K$ decay at infinity at least with the rate $\frac{1}{1+|x|^{2}}$ by Theorem 4.1. And they possess at least Hölder continuity in any bounded region due to classical elliptic estimates for the linear Stokes operator.
We now try to solve a fixed-point problem.
(4.3) ${\boldsymbol{v}}=K({\boldsymbol{v}},\sigma)\quad\text{in }X$
that satisfies the following:
  • (1)
    $K$ is continuous; $K(\cdot,\sigma)$ is compact for each $\sigma$.
  • (2)
    It is uniquely solvable in $X$ for small $\sigma$ by classical result; see for example [10] and [8, Proposition 3]
  • (3)
    We have a priori estimates in $X$ for solutions $({\boldsymbol{v}},\sigma)\in X\times[0,1]$ by Theorem 4.1.
By the Leray-Schauder theorem, there is a solution ${\boldsymbol{v}}\in X$ of (4.3) with $\sigma=1$. It follows that $u_{\textrm{re}}=E{\boldsymbol{v}}$ satisfies the nonhomogeneous Stokes system (3.13) with $f=-(e^{t\Delta}u_{0}+u_{\textrm{re}})\otimes(e^{t\Delta}u_{0}+u_{\textrm{re}})$, and hence $u=e^{t\Delta}u_{0}+u_{\textrm{re}}$ is a self-similar solution of (1.1) with initial data $u_{0}$. It is easy to check the solution we find satisfies the Definition 1.1. Next, we prove the pointwise estimates in Theorem 1.1. We have by Theorem 4.1 and Lemma 2.1 that
$|u(x,t)|=\frac{1}{\sqrt{t}}\left|v\left(\frac{x}{\sqrt{t}}\right)\right|\leq C(A,\beta)\frac{1}{\sqrt{t}}\frac{1}{1+\frac{|x|}{\sqrt{t}}}=\frac{C(A,\beta)}{|x|+\sqrt{t}},$
and
$|u(x,t)-e^{t\Delta}u_{0}|=\frac{1}{\sqrt{t}}\left|v_{\textrm{re}}\left(\frac{x}{\sqrt{t}}\right)\right|\leq C(A,\beta)\frac{1}{\sqrt{t}}\frac{1}{1+\frac{|x|^{2}}{t}}=\frac{C(A,\beta)\sqrt{t}}{|x|^{2}+t},$
and
$|\nabla u(x,t)-\nabla e^{t\Delta}u_{0}|$ $=\frac{1}{t}\left|\nabla v_{\textrm{re}}\left(\frac{x}{\sqrt{t}}\right)\right|\leq C(A,\beta)\frac{1}{t}\frac{1}{1+\frac{|x|^{2+\beta}}{t^{\frac{2+\beta}{2}}}}$
$\leq\frac{C(A,\beta)t^{\frac{\beta}{2}}}{(|x|+\sqrt{t})^{2+\beta}}.$
At last, if the initial data $u_{0}\in C^{2}(S^{1})$, we have the following optimal decay rate estimates
$|\nabla u(x,t)-\nabla e^{t\Delta}u_{0}|$ $=\frac{1}{\sqrt{t}}\left|\nabla v_{\textrm{re}}\left(\frac{x}{\sqrt{t}}\right)\right|$
$\leq C(\|u_{0}\|_{C^{2}(S^{1})})\frac{t}{(|x|+\sqrt{t})^{3}}\ln\left(1+\frac{|x|}{\sqrt{t}}\right).$
Appendix A Proof of Lemma 2.1
In this appendix, we prove Lemma 2.1. Before we proceed, we give the following proposition concerning the difference estimates for (1)(-1)-homogeneous, locally Hölder continuous vector field, which is needed in the proof of Lemma 2.1.
Proposition A.1.
Let $u_{0}(x)$ be a (1)(-1)-homogeneous vector field in $\mathbb{R}^{n}\setminus\{0\}$ ($n\geq 2$) such that its restriction to the unit sphere satisfies $u_{0}\in C^{0,\beta}(S^{n-1})$ with $0<\beta<1$. Let $A=\|u_{0}\|_{C^{0,\beta}(S^{n-1})}$. Then for any $y,z\in\mathbb{R}^{n}\setminus\{0\}$, the following estimate holds:
(A.1) $|u_{0}(z)-u_{0}(y)|\leq C(A)\frac{|z-y|^{\beta}}{\min(|z|,|y|)^{1+\beta}}.$
Proof.
Let $y,z\in\mathbb{R}^{n}\setminus\{0\}$. We use polar coordinates $y=\rho\theta$ and $z=r\omega$, where $\rho=|y|$, $r=|z|$, and $\theta,\omega\in S^{n-1}$. By the (1)(-1)-homogeneity of $u_{0}$, we have $u_{0}(y)=\rho^{-1}u_{0}(\theta)$ and $u_{0}(z)=r^{-1}u_{0}(\omega)$. The difference is decomposed as:
$|u_{0}(z)-u_{0}(y)|$ $=\left|\frac{1}{r}u_{0}(\omega)-\frac{1}{\rho}u_{0}(\theta)\right|$
$=\left|\frac{1}{r}(u_{0}(\omega)-u_{0}(\theta))+\left(\frac{1}{r}-\frac{1}{\rho}\right)u_{0}(\theta)\right|$
$\leq\frac{1}{r}|u_{0}(\omega)-u_{0}(\theta)|+|u_{0}(\theta)|\left|\frac{1}{r}-\frac{1}{\rho}\right|=:\text{Term I}+\text{Term II}.$
Step 1: Estimating Term I (Angular Term). Since $u_{0}\in C^{\beta}(S^{n-1})$, we have $|u_{0}(\omega)-u_{0}(\theta)|\leq A|\omega-\theta|^{\beta}$. Utilizing the geometric inequality
$\left|\omega-\theta\right|=\left|\frac{z}{|z|}-\frac{y}{|y|}\right|\leq\frac{2|z-y|}{\max(|z|,|y|)},$
we obtain
$\frac{1}{r}|u_{0}(\omega)-u_{0}(\theta)|\leq\frac{A}{r}\left(\frac{2|z-y|}{\max(r,\rho)}\right)^{\beta}\leq\frac{C(A)}{\min(r,\rho)^{1+\beta}}|z-y|^{\beta}.$
Step 2: Estimating Term II (Radial Term). Without loss of generality, assume $\rho\leq r$. We prove Term II is bounded by $C\rho^{-(1+\beta)}|z-y|^{\beta}$ by considering two cases for $|r-\rho|$:
Case A: Near-field radial difference ($|r-\rho|\leq\frac{\rho}{2}$). In this case, $r\geq\rho$ and $|r-\rho|\leq|z-y|$. We have:
$\left|\frac{1}{r}-\frac{1}{\rho}\right|=\frac{|r-\rho|}{r\rho}=\frac{|r-\rho|^{\beta}\cdot|r-\rho|^{1-\beta}}{r\rho}\leq\frac{|z-y|^{\beta}\cdot(\rho/2)^{1-\beta}}{\rho^{2}}=\frac{1}{2^{1-\beta}}\frac{|z-y|^{\beta}}{\rho^{1+\beta}}.$
Case B: Far-field radial difference ($|r-\rho|>\frac{\rho}{2}$). In this case, $|z-y|\geq|r-\rho|>\frac{\rho}{2}$, which implies $1<\frac{2^{\beta}|z-y|^{\beta}}{\rho^{\beta}}$. Since $r\geq\rho$, we have:
$\left|\frac{1}{r}-\frac{1}{\rho}\right|\leq\frac{1}{r}+\frac{1}{\rho}\leq\frac{2}{\rho}<\frac{2}{\rho}\cdot\frac{2^{\beta}|z-y|^{\beta}}{\rho^{\beta}}=\frac{2^{1+\beta}|z-y|^{\beta}}{\rho^{1+\beta}}.$
Combining both cases, we have $\text{Term II}\leq\frac{CA}{\min(r,\rho)^{1+\beta}}|z-y|^{\beta}$. Summing the estimates for Term I and Term II yields (A.1). ∎
We can now prove Lemma 2.1.
Proof of Lemma 2.1.
The smoothness of $v_{0}$ follows directly from the classical regularity theory for the heat equation. We now proceed to prove the estimate (2.12). Throughout the proof, $C(A)$ ($C(A,\beta)$) denotes a constant depending only on AA (AA and $\beta$), which may vary from line to line. We denote $G(x)=\frac{1}{4\pi}e^{-|x|^{2}/4}$ as the heat kernel at $t=1$ on $\mathbb{R}^{2}$ in this proof. Then
$v_{0}(y)=\int_{\mathbb{R}^{2}}G(y-z)u_{0}(z)dz,$
and by our assumption $u_{0}(z)=\frac{1}{|z|}u_{0}\left(\frac{z}{|z|}\right)$ with $\|u_{0}\|_{C^{0,\beta}(S^{1})}\leq A$.
The case $k=0$. For the region $|y|\leq 1$, $v_{0}(y)$ is clearly bounded. Specifically, when $|y|\leq 1$,
(A.2) $|v_{0}(y)|$ $\leq\left(\int_{|z|\leq 2}+\int_{|z|\geq 2}\right)e^{-\frac{|y-z|^{2}}{4}}\frac{C(A)}{|z|}dz$
$\leq\left(\int_{|z|\leq 2}\frac{C(A)}{|z|}dz+\int_{|z|\geq 2}e^{-\frac{|y-z|^{2}}{4}}\frac{C(A)}{2}dz\right)$
$\leq C(A)+C(A)\int_{|z|\geq 2}e^{-\frac{|z|^{2}}{16}}dz\leq C(A).$
For $|y|>1$,
$|v_{0}(y)|$ $\leq\left(\int_{|z|\leq\frac{|y|}{2}}+\int_{\frac{|y|}{2}\leq|z|\leq 2|y|}+\int_{|z|\geq 2|y|}\right)e^{-\frac{|y-z|^{2}}{4}}\frac{C(A)}{|z|}dz$
$=I_{1}+I_{2}+I_{3}.$
When $|z|\leq\frac{|y|}{2}$, we know $|y-z|\geq\frac{|y|}{2}$, and then
$I_{1}\leq e^{-\frac{|y|^{2}}{16}}\int_{|z|\leq\frac{|y|}{2}}\frac{C(A)}{|z|}dz\leq C(A)e^{-\frac{|y|^{2}}{16}}|y|.$
For $I_{2}$, we first note that $\left\{\frac{|y|}{2}\leq|z|\leq 2|y|\right\}\subset\left\{z:|y-z|\leq 3|y|\right\}$. Hence,
$I_{2}$ $\leq\int_{\left\{z:|y-z|\leq 3|y|\right\}}e^{-\frac{|y-z|^{2}}{4}}\frac{C(A)}{|y|}dz$
$\leq\frac{C(A)}{|y|}\int_{\left\{|\tilde{z}|\leq 3|y|\right\}}e^{-\frac{|\tilde{z}|^{2}}{4}}dz\leq\frac{C(A)}{|y|}.$
For $I_{3}$, we note that when $|z|\geq 2|y|$, we have $|y|\leq\frac{|z|}{2}$, $|y-z|\geq\frac{|z|}{2}$ and also $|y-z|\geq|y|$. Therefore, we can estimate $I_{3}$ as
$I_{3}$ $\leq\frac{C(A)}{|y|}\int_{|z|\geq 2|y|}e^{-\frac{|y-z|^{2}}{8}}e^{-\frac{|y-z|^{2}}{8}}dz$
$\leq\frac{C(A)}{|y|}e^{-\frac{|y|^{2}}{8}}\int_{|z|\geq 2|y|}e^{-\frac{|z|^{2}}{32}}dz\leq\frac{C(A)}{|y|}e^{-\frac{|y|^{2}}{8}}.$
This shows that for $|y|>1$, we have
(A.3) $|v_{0}(y)|\leq\frac{C(A)}{|y|}.$
Combining (A.2) and (A.3), we obtain
$|v_{0}(y)|\leq\frac{C(A)}{1+|y|}.$
This completes the proof of the (2.12) for $k=0$.
The case $k=1$. Next, we prove (2.12) for $k=1$. We consider first the component $\beta=1$, and we have that
$\partial_{i}v_{0}(y)=\frac{1}{4\pi}\int_{\mathbb{R}^{2}}e^{-\frac{|y-z|^{2}}{4}}\left(-\frac{1}{2}\right)(y_{i}-z_{i})u_{0}(z)dz,~i=1,2.$
We have
$|\partial_{i}v_{0}(y)|$ $\leq\int_{\mathbb{R}^{2}}e^{-\frac{|y-z|^{2}}{4}}|y_{i}-z_{i}|\frac{A}{|z|}dz$
$\leq\int_{\mathbb{R}^{2}}e^{-\frac{|y-z|^{2}}{4}}|y-z|\frac{A}{|z|}dz.$
Hence, when $|y|\leq 1$,
$|\partial_{i}v_{0}(y)|$ $\leq\left(\int_{|z|\leq 2}+\int_{|z|\geq 2}\right)e^{-\frac{|y-z|^{2}}{4}}|y-z|\frac{C(A)}{|z|}dz$
$\leq\left(\int_{|z|\leq 2}\frac{C(A)}{|z|}dz+\int_{|z|\geq 2}e^{-\frac{|y-z|^{2}}{4}}|y-z|C(A)dz\right)$
$\leq C(A)+C(A)\int_{|z|\geq 2}e^{-\frac{|z|^{2}}{16}}dz\leq C(A).$
On the other hand, when $|y|\geq 1$, we rewrite
$\partial_{i}v_{0}(y)$ $=\frac{1}{4\pi}\int_{|z|\leq\frac{1}{2}}+\int_{|z|\geq\frac{1}{2}}e^{-\frac{|y-z|^{2}}{4}}\left(-\frac{1}{2}\right)(y_{i}-z_{i})u_{0}(z)dz$
$=\frac{1}{4\pi}\int_{|z|\leq\frac{1}{2}}e^{-\frac{|y-z|^{2}}{4}}\left(-\frac{1}{2}\right)(y_{i}-z_{i})u_{0}(z)dz+\frac{1}{4\pi}\int_{|z|\geq\frac{1}{2}}e^{-\frac{|y-z|^{2}}{4}}\partial_{i}u_{0}(z)dz$
$\quad-\frac{1}{4\pi}\int_{\partial B_{\frac{1}{2}}}e^{-\frac{|y-z|^{2}}{4}}u_{0}(z)\cdot nd\sigma_{z}=J_{1}+J_{2}+J_{3}.$
Similar to estimate of $v_{0}$ above, it is easy to find that for $|y|>1$.
$|J_{1}|\leq\int_{|z|\leq\frac{1}{2}}e^{-\frac{|y-z|^{2}}{4}}|y-z|\frac{C(A)}{|z|}dz\leq\int_{|z|\leq\frac{1}{2}}e^{-\frac{|y|^{2}}{16}}\frac{|y|}{2}\frac{C(A)}{|z|}dz\leq\frac{C(A)}{1+|y|^{2}}.$
$|J_{2}|\leq\int_{|z|\geq\frac{1}{2}}e^{-\frac{|y-z|^{2}}{4}}\frac{C(A)}{|z|^{2}}dz\leq\frac{C(A)}{1+|y|^{2}}.$
$|J_{3}|\leq C(A)\int_{\partial B_{\frac{1}{2}}}e^{-\frac{|y-z|^{2}}{4}}d\sigma_{z}\leq\frac{C(A)}{1+|y|^{2}}.$
We finish the proof of (2.12) for $k=1$ and $\beta=1$.
Next, we consider the case for $k=1$ and $0<\beta<1$. The derivative $\partial_{i}v_{0}(y)$ can be expressed as:
$\partial_{i}v_{0}(y)=\int_{\mathbb{R}^{2}}\partial_{y_{i}}G(y-z)u_{0}(z)\,dz,$
where $G(x)=\frac{1}{4\pi}e^{-|x|^{2}/4}$ is the heat kernel at $t=1$. Since $\int_{\mathbb{R}^{2}}\partial_{y_{i}}G(y-z)\,dz=0$, we can utilize the following identity to handle the singularity and the decay:
(A.4) $\partial_{i}v_{0}(y)=\int_{\mathbb{R}^{2}}\partial_{y_{i}}G(y-z)[u_{0}(z)-u_{0}(y)]\,dz.$
The proof of boundedness of $\partial_{i}v_{0}$ when $|y|\leq 1$ is similar to the above computations. For $|y|>1$, we divide the integration domain into three regions similar as above: $L_{1}=\{z:|z|\leq\frac{|y|}{2}\}$ and $L_{2}=\{z:\frac{|y|}{2}\leq|z|\leq 2|y|\}$ and $L_{3}=\{z:|z|\geq 2|y|\}$.
Case 1: Region $L_{1}$. In region $L_{1}$, we revert to the original form of the integral using the identity (A.4):
$\int_{L_{1}}\partial_{y_{i}}G(y-z)u_{0}(z)\,dz-u_{0}(y)\int_{L_{1}}\partial_{y_{i}}G(y-z)\,dz=:M_{1}+M_{2}.$
For $M_{1}$, we have
$|M_{1}|\leq C(A)e^{-\frac{|y|^{2}}{16}}\int_{|z|\leq|y|/2}\frac{1}{|z|}\,dz\leq C(A)e^{-\frac{|y|^{2}}{16}}|y|.$
For $M_{2}$, since $|u_{0}(y)|\leq C|y|^{-1}$, we have:
$|M_{2}|\leq C(A)e^{-\frac{|y|^{2}}{16}}\frac{1}{|y|}\int_{|z|\leq|y|/2}\,dz\leq C(A)e^{-\frac{|y|^{2}}{16}}|y|.$
Case 2: Region $L_{2}$. It follows from Proposition A.1 in region $L_{2}$ that
$|u_{0}(z)-u_{0}(y)|\leq C(A)\frac{|z-y|^{\beta}}{\min(|z|,|y|)^{1+\beta}}\leq\frac{C(A)}{|y|^{1+\beta}}|z-y|^{\beta}.$
In this region, for $z\in L_{2}$, we have $|y-z|\leq 3|y|$. We get
$\left|\int_{L_{2}}\partial_{y_{i}}G(y-z)[u_{0}(z)-u_{0}(y)]\,dz\right|$ $\leq\frac{C(A)}{|y|^{1+\beta}}\int_{|\tilde{z}|\leq 3|y|}|\partial_{i}G(\tilde{z})||\tilde{z}|^{\beta}\,d\tilde{z}$
$\leq\frac{C(A,\beta)}{|y|^{1+\beta}}.$
Case 3: Region $L_{3}$. The estimates in the region are similar to Region $L_{1}$. We conclude that
$\left|\int_{L_{3}}\partial_{y_{i}}G(y-z)u_{0}(z)\,dz-u_{0}(y)\int_{L_{3}}\partial_{y_{i}}G(y-z)\,dz\right|\leq C(A)e^{-\frac{|y|^{2}}{16}}$
Combining the estimates for $L_{1}$, $L_{2}$ and $L_{3}$, we obtain that for $\beta\in(0,1)$:
$|\partial_{i}v_{0}(y)|\leq\frac{C(A,\beta)}{1+|y|^{1+\beta}}.$
The proofs for $k\geq 2$ follow by applying similar arguments to the higher-order derivatives of the heat kernel, and we finish the proof. ∎
Acknowledgement. The research of Gui is supported by NSFC Key Program (Grant No. 12531010), University of Macau research grants CPG2024-00016-FST, CPG2025-00032-FST, CPG2026-00027-FST, SRG2023-00011-FST, MYRG-GRG2023-00139-FST-UMDF, UMDF Professorial Fellowship of Mathematics, Macao SAR FDCT 0003/2023/RIA1 and Macao SAR FDCT 0024/2023/RIB1. The research of Xie is partially supported by NSFC grants 12571238 and 12426203.